### v1

# Taylor-Fourier integration

Abstract
In this paper we introduce an algorithm which provides approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$. The proposed approximate solutions are written in closed form as functions $X(t,\omega\,t)$ where $X(t,\theta)$ is, (i) a truncated Fourier series in $\theta$ for fixed $t$, and (ii) a truncated Taylor series in $t$ for fixed $\theta$ (that is the reason for the name of the proposed integrators). Such approximations are intended to be uniformly accurate in $\omega$ (in the sense that their accuracy is not deteriorated as $\omega\to\infty$). This feature implies that Taylor-Fourier approximations become more efficient than the application of standard numerical integrators for sufficiently high basic frequency $\omega$. The main goal of the paper is to propose a procedure to efficiently compute such approximations by combining power series arithmetic techniques and the FFT algorithm. We present numerical experiments that demonstrate the effectiveness of our approximation method through its application to well-known problems of interest.
Keywords: Taylor series, Fourier methods, oscillatory problems, numerical integrators.
1 Introduction
In this paper we consider differential systems of the form
$\frac{d}{dt}{\bf x}=\omega A{\bf x}+{\bf g}({\bf x}),$ (1)
where ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a smooth map, $\omega$ is a positive real parameter, and $A$ is a real $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$. Hence, the matrix exponential $e^{\theta A}$ is $(2\pi)$-periodic in $\theta$ and, therefore, the solution ${\bf x}(t)=e^{t\omega A}{\bf x}(0)$ of (1) with ${\bf g}\equiv{\bf 0}$ is $(2\pi/\omega)$-periodic in $t$. If $\omega$ is large, the solution of the homogeneous problem associated to (1) is highly oscillatory. Examples of differential systems (1) with these properties can be found, for instance, when dealing with semi-discretized dispersive partial differential equations [3, 9, 17], or in some formulations of the equations of motion in satellite dynamics [16].
We focus on approximating the solution ${\bf x}(t)$ of (1) supplemented with the initial condition
${\bf x}(0)={\bf x}_{0}\in\mathbb{R}^{D}.$ (2)
With the change of variables
${\bf x}(t)=e^{t\omega A}{\bf y}(t),$ (3)
the initial value problem (1)–(2) reads
$\frac{d}{dt}{\bf y}={\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf x}_{0},$ (4)
with
${\bf f}(\theta,{\bf y})=e^{-\theta A}{\bf g}(e^{\theta A}{\bf y}),$ (5)
which is $(2\pi)$-periodic in $\theta$.
Provided that the basic frequency $\omega$ is high enough, solving numerically (4) by a Runge-Kutta method demonstrates superior efficiency compared to the direct application of the same Runge-Kutta scheme to the original equation (1).
In the context of problems of the form (1) obtained from the space semi-discretization of dispersive PDEs, the numerical integration with a Runge-Kutta method of the transformed system (4), as a means to solve (1)–(2), leads to the so-called Lawson’s generalized Runge-Kutta methods [11, 9].
A completely different context where problems of the considered form arise is the numerical orbit propagation of artificial satellites. The equations of motion of the satellite can be brought into the form (1), for instance, by using the Kustaanheimo-Stiefel (KS) formulation [10]. In this context, (4) is referred to as the VOP (variation of parameters) formulation of the system (1). In particular, a VOP formulation of the KS equations was presented and analyzed by Stiefel and Scheifele in [16]. Other VOP formulations (obtained by using different sets of coordinates) of the equations of motion in satellite dynamics are also of the form (4).
From now on we will focus on computing approximations to the solution ${\bf y}(t)$ of the initial value problem (4) which are valid on an appropriate interval around $t=0$. In the highly oscillatory case, for high enough basic frequency $\omega$, the larger is the basic frequency $\omega$, the more computational effort will be required by the Runge-Kutta method to achieve a prescribed accuracy. In the present work, we aim at obtaining a closed form approximation to the solution ${\bf y}(t)$ of (4) that exhibits uniform accuracy across varying frequencies $\omega$. This implies that comparable accuracy is attained with the same computational effort even for arbitrarily high frequencies. The Taylor-Fourier approximations that we propose here draw inspiration from previous works on the numerical integration of oscillatory problems (see e.g. [4], [7], and references therein), and are related to the methods proposed in [5] and in [2].
For a prescribed positive integer $M$, we propose to compute a sequence
$\left\{{\bf y}^{[1]}(t),{\bf y}^{[2]}(t),\ldots,{\bf y}^{[d]}(t),\ldots\right\}$
of approximations to the solution ${\bf y}(t)$ of (4) of the form
${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right),\quad d=1,2,\ldots,$ (6)
with appropriate coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, and starting from ${\bf y}^{[0]}(t)\equiv{\bf x}_{0}$ (this means that we can use (6) with $d=0$, setting ${\bf y}_{0,0}^{[0]}={\bf x}_{0}$). We will refer to (6) as the (M,d)-Taylor-Fourier approximation to the solution of (4).
In order to compute the successive approximations (6) to the solution of (4), for $d=0,1,\ldots$, we do the following:
  • Given ${\bf y}^{[d]}(t)$, we first insert it into the right hand side of (4), and then we determine an approximation ${\bf z}^{[d]}(t)$ to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ of the form
    ${\bf z}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf z}_{k,j}^{[d]}\right).$ (7)
    This approximation, to be defined precisely in Section 2, is determined by combining truncated Taylor expansions and trigonometric interpolation of periodic functions. That is the reason for the name of the proposed integrators.
  • Once ${\bf z}^{[d]}(t)$ has been obtained, we use it as the right-hand side of (4) and integrate to obtain
[…]
    It is straightforward to check that, by construction, ${\bf y}_{k,d+1}^{[d+1]}=0$ for $k\neq 0$.
The aim of the present paper is: (i) to introduce and give a precise definition of Taylor-Fourier approximations, (ii) to propose a procedure to compute them efficiently by making extensive use of the FFT algorithm, and (iii) to illustrate the application of the proposed approximation method to certain well known problems of interest. Our findings suggest that, provided that the basic frequency $\omega$ is high enough, Taylor-Fourier approximations can achieve comparable precision to standard numerical integration methods applied to the transformed system (4), while demanding less CPU time.
We want to stress that employing Taylor-Fourier approximations may be of interest even in cases where the application of standard numerical integrators to the transformed system (4) proves to be very efficient. Indeed, the algorithm introduced in the present paper provides, for given $d$ and $M$, closed expressions
${\bf X}(t)=e^{t\omega A}\left(\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right)$ (9)
of approximate solutions of (1)-(2), valid for a relatively large time interval and that, once computed, can be efficiently evaluated at arbitrary times $t$. This is in contrast with most time integrators commonly used, which generate approximations to the exact solution of the problem either at a prescribed final time $T$ or at a finite set $0<t_{1}<t_{2}<\cdots\leq T$ of intermediate times. (In the later case, approximations to the solutions at arbitrary times $t\in[0,T]$ are typically obtained with piecewise interpolating polynomials.)
In our numerical experiments, we will compare the $(M,d)$-Taylor-Fourier approximations (6) to the solution ${\bf y}(t)$ of (4) (equivalently, the approximation (9) to the solution ${\bf x}(t)$ of (1)–(2)) with the approximations obtained by numerically integrating (4) with highly efficient Runge-Kutta methods (equivalently, numerically integrating (1)–(2) with Lawson’s generalized Runge-Kutta methods). A state-of-the-art highly performant library for the numerical integration of differential equations, DifferentialEquation.jl [15], implemented in julia language [1], is freely available and can be used to integrate the transformed system (4) with optimized adaptive explicit Runge-Kutta integrators. In order to compare such numerical solvers with the integrators proposed in this paper, we have carried out a careful implementation of the Taylor-Fourier methods in julia language, paying special attention to computational aspects such as memory management and efficient power series arithmetic.
In Section 2 we give a precise definition of $(M,d)$-Taylor-Fourier approximations, and demonstrate their applicability to problem classes extending beyond real semi-linear problems of the form (1). Section 3 is devoted to describe an algorithm designed to efficiently compute Taylor-Fourier approximations. In Section 4 we present a couple of relevant numerical examples where the performance of Taylor-Fourier integrators is analyzed. Finally, in Section 5, we summarize our conclusions and outline avenues for future research on Taylor-Fourier methods.
2 Taylor-Fourier methods
2.1 Definition of Taylor-Fourier approximations
So far, we have not specified how to determine the approximation (7) to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ in the procedure introduced in Section 1. We first note that (6) and (7) can be rewritten as
${\bf y}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\omega t),\qquad{\bf z}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\omega t),$
respectively, where each component of the $D$-vector functions ${\bf Y}_{j}^{[d]}(\theta)$ and ${\bf Z}_{j}^{[d]}(\theta)$ is a trigonometric polynomial of degree $M$ in the angular variable $\theta$, that is,
${\bf Y}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf y}_{k,j}^{[d]},\qquad{\bf Z}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j}^{[d]},$
with coefficients ${\bf y}_{k,j}^{[d]},{\bf z}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$.
Given a positive integer $M$ and ${\bf Y}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$, we will determine ${\bf Z}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$ by imposing the following conditions:
  • For each $\theta_{n}=n\pi/M$, $n=0,1,2,\ldots,2M-1$,
    ${\bf f}\left(\theta_{n},\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta_{n})\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta_{n})+\mathcal{O}(t^{d+1}){\color[rgb]{1,0,0},}$ (10)
    as $t\to 0$.
  • For $j=0,1,\ldots,d$,
    ${\bf z}_{-M,j}^{[d]}={\bf z}_{M,j}^{[d]}.$ (11)
The first condition is aimed at ensuring that
$\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta)\approx{\bf f}\left(\theta,\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)\right)$
for all $\theta\in[0,2\pi]$ and for sufficiently small $t\in\mathbb{R}$, whereas the second one is included to ensure uniqueness.
To sum up, given a positive integer $M$, the successive approximations (6) to the solution of (4) are uniquely determined by (10)–(11) and (8). We will refer to the approximation (6) as the $(M,d)$-Taylor-Fourier approximation.
2.2 Application to more general equations
The definition of Taylor-Fourier approximations can be extended without modifications to initial value problems of the form (4) in the complex domain. In particular, Taylor-Fourier approximations can be successfully applied to equations of the form (1) defined in the complex domain, provided that $A$ is a complex $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$, and the map ${\bf g}:\mathbb{C}^{D}\to\mathbb{C}^{D}$ is such that ${\bf g}({\bf x})={\bf G}({\bf x},{\bf\bar{x}})$, where ${\bf G}:\mathbb{C}^{2D}\to\mathbb{C}^{D}$ is a polynomial map.
Going back to equations in the real domain, Taylor-Fourier methods are not limited to semi-linear equations of the form (1). They can be applied more generally to problems of the form
$\frac{d}{dt}{\bf x}=\omega\,{\bf r}({\bf x})+{\bf g}({\bf x}),\quad{\bf x}(0)={\bf x}_{0},$ (12)
where ${\bf r}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ and ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ are smooth maps and $\omega$ is a positive real parameter, provided that the following two assumptions hold:
  1. 1.
    For each $t\in\mathbb{R}$ there exists a $t$-flow map $\varphi_{t}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ of the differential system $\frac{d}{dt}{\bf x}={\bf r}({\bf x})$. That is, for each $t\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{D}$,
    $\frac{\partial}{\partial t}\varphi_{t}({\bf x})={\bf r}(\varphi_{t}({\bf x})){\color[rgb]{1,0,0}.}$ (13)
    In that case, problem (12) is transformed with the change of variables
    ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$
    into (4), where
    ${\bf f}(\theta,{\bf y})=\varphi^{\prime}_{\theta}({\bf y})^{-1}{\bf g}(\varphi_{\theta}({\bf y})).$ (14)
    Here, $\varphi^{\prime}_{\theta}({\bf y})$ denotes the Jacobian matrix of the map $\varphi_{\theta}$ evaluated at the vector ${\bf y}$.
  2. 2.
    The function ${\bf f}(\theta,{\bf y})$ defined in (14) is $(2\pi)$-periodic in $\theta$.
Under these two assumptions, the solution ${\bf x}(t)$ of (12) can be approximated by
${\bf x}^{[d]}(t)=\varphi_{\omega\,t}\left({\bf y}^{[d]}(t)\right),$ (15)
where
${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}$
is the $(M,d)$-Taylor-Fourier approximation to the solution of (4)–(14).
In particular, the above-mentioned two assumptions hold for systems of the form
$\begin{split}\frac{d}{dt}\phi_{j}&=k_{j}\omega+g_{j}({\bf x}),\quad j=1,\ldots,d,\\ \frac{d}{dt}a_{j}&=g_{d+j}({\bf x}),\quad j=1,\ldots,D-d,\\ \end{split}$ (16)
where ${\bf x}=(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})$, $k_{j}\in\mathbb{Z}$, and for each $j$, $g_{j}$ is $2\pi$-periodic in each of the angle variables $\phi_{j}$. In that case,
$\varphi_{t}(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})=\left(\begin{matrix}\phi_{1}+k_{1}\,t\\ \vdots\\ \phi_{d}+k_{d}\,t\\ a_{1}\\ \vdots\\ a_{D-d}\\ \end{matrix}\right),$
and the change of variables ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$, with ${\bf y}=(\tilde{\phi}_{1},\ldots,\tilde{\phi}_{d},\tilde{a}_{1},\ldots,\tilde{a}_{D-d})$, transforms the system (16) into
$\frac{d}{dt}\tilde{\phi}_{j}$ $=g_{j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,d,$
$\frac{d}{dt}\tilde{a}_{j}$ $=g_{d+j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,D-d.$
3 Efficient implementation of Taylor-Fourier integrators
We will next show, for given positive integers $M$ and $d$, and given vector coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$, how to compute efficiently the new coefficients ${\bf y}_{k,j}^{[d+1]}\in\mathbb{C}^{D}$, $10\leq j\leq d+$ 1, $-M\leq k\leq M$. This computation is carried out in four sequential steps (for simplicity of notation, we omit superscripts $[d]$ in ${\bf Y}_{j}^{[d]}$ and ${\bf Z}_{j}^{[d]}$, when they are not necessary). We recall that $\theta_{n}=n\pi/M,\quad n=0,1,\ldots,2M-1$.
3.1 Step 1: Evaluation of trigonometric polynomials
For $n=0,1,\ldots,2M-1$, we calculate the vector coefficients
${\bf Y}_{n,j}={\bf Y}_{j}(\theta_{n})=\sum_{k=-M}^{M}e^{ikn\pi/M}{\bf y}_{k,j}^{[d]},\quad j=0,1,\ldots,d,$ (17)
by applying the inverse DFT
${\bf Y}_{n,j}=\frac{1}{2M}\sum_{k=0}^{2M-1}e^{ikn\pi/M}\hat{\bf Y}_{k,j},\quad n=0,1,\ldots,2M-1,$
where for $j=0,1,\ldots,d$,
$\begin{array}[]{l}\hat{\bf Y}_{0,j}=2M\,{\bf y}_{0,j}^{[d]},\quad\hat{\bf Y}_{M,J}=2M({\bf y}_{M,j}^{[d]}+{\bf y}_{-M,j}^{[d]}),\\[8.0pt] \hat{\bf Y}_{k,j}=2M\,{\bf y}_{k,j}^{[d]},\quad\hat{\bf Y}_{2M-k,j}=2M{\bf y}_{-k,j}^{[d]},\quad k=1,\ldots,M-1.\end{array}$
3.2 Step 2: Truncated Taylor expansions
For each $n=0,1,\ldots,2M-1$, we compute ${\bf Z}_{n,j}$, $j=0,1,\ldots,d$, such that
${\bf f}\left({\theta_{n},}\sum_{j=0}^{d}t^{j}\,{\bf Y}_{n,j}\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}+{\cal O}(t^{d+1}).$
This means that we obtain the Taylor polynomial of degree $d$ of the left hand side of the previous equality. The computation of the required Taylor expansion can be done in a very efficient way when appropriate tools to perform power series arithmetic are available.
3.3 Step 3: Trigonometric interpolation
For each $j=0,1,\ldots,d$, we determine the coefficients ${\bf z}_{k,j}$, $-M\leq k\leq M$, of the trigonometric polynomial
${\bf Z}_{j}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j},$
that satisfy the interpolatory conditions
${\bf Z}_{j}(\theta_{n})={\bf Z}_{n,j},\quad n=0,1,2,\ldots,2M-1,$
together with ${\bf z}_{-M,j}={\bf z}_{M,j}$.
This task can be efficiently done by computing the discrete Fourier Transform (DFT)
$\hat{\bf Z}_{k,j}=\sum_{n=0}^{2M-1}e^{-ikn\pi/M}\,{\bf Z}_{n,j},\quad k=0,1,\ldots,2M-1,$ (18)
and setting
$\begin{array}[]{l}{\bf z}_{0,j}=\dfrac{1}{2M}\hat{\bf Z}_{0,j},\quad{\bf z}_{M,j}={\bf z}_{-M,j}=\dfrac{1}{4M}\hat{\bf Z}_{M,J},\\[8.0pt] {\bf z}_{k,j}=\dfrac{1}{2M}\hat{\bf Z}_{k,j},\quad{\bf z}_{-k,j}=\dfrac{1}{2M}\hat{\bf Z}_{2M-k,j},\quad k=1,\ldots,M-1.\end{array}$ (19)
Remark: For systems of the form (4) in the real domain, the symmetries present in the complex vector resulting from the application of DFT to a real vector can be used to compute it more efficiently (with the real FFT algorithm) than in the general case of complex vectors. In particular, this implies that each ${\bf z}_{-k,j}$ coincides with the complex conjugate of ${\bf z}_{k,j}$, which allows us to halve the number of coefficients that must be computed and saved.
3.4 Step 4: Computing quadratures
Given
${\bf z}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf z}_{k,j}\right),$
we want to write
[…]
To simplify the notation, we will drop the superscript $[d+1]$ from the vector coefficients ${\bf y}_{k,j}^{[d+1]}$ in the remaining of the present subsection, so that we will write
${\bf y}^{[d+1]}(t)=t^{d+1}\,{\bf y}_{0,d+1}+\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}\right).$
First, we notice that ${\bf z}(t)$ can be written as
${\bf z}(t)={\bf p}(t)+\sum_{j=0}^{d}t^{j}\,\tilde{{\bf Z}}_{j}(\omega t),$ (21)
where
${\bf p}(t)=\sum_{j=0}^{d}t^{j}\,{\bf z}_{0,j},\quad\tilde{{\bf Z}}_{j}(\theta)=\sum_{k=1}^{M}(e^{ik\theta}{\bf z}_{k,j}+e^{-ik\theta}{\bf z}_{-k,j}),\quad j=0,\ldots,d.$ (22)
Therefore,
${\bf y}(t)={\bf y}_{0}+{\bf q}(t)+\sum_{j=0}^{d}t^{j}\,\tilde{{\bf Y}}_{j}(\omega t),$ (23)
with
$\tilde{{\bf Y}}_{j}(\theta)$ $=$ $\sum_{k=1}^{M}(e^{ik\theta}{\bf y}_{k,j}+e^{-ik\theta}{\bf y}_{-k,j}),\quad j=0,\ldots,d,$ (24)
[…]
The coefficients ${\bf y}_{k,j},{\bf y}_{-k,j}$, $j=0,1,\ldots,{d+1}$, $k=1,\ldots,M$, can be efficiently obtained as follows:
  • Set $\tilde{{\bf Y}}_{d+1}(\theta)\equiv{\bf 0}$, and notice that taking derivatives in (23) with respect to $\theta=\omega t$, and using (20) and (21) one gets
    $\omega\frac{d}{d\theta}\tilde{{\bf Y}}_{j}(\theta)=\tilde{{\bf Z}}_{j}(\theta)-(j+1)\tilde{{\bf Y}}_{j+1}(\theta),\quad j=0,1,\ldots,d.$
    This means that setting ${\bf y}_{k,d+1}={\bf 0}$ for $k\neq 0$, we can compute recursively for $j=d,d-1,\ldots,1,0$,
    ${\bf y}_{k,j}=\frac{1}{ik\omega}\left({\bf z}_{k,j}-(j+1){\bf y}_{k,j+1}\right).$
  • For $k=0$, one gets from (25) and (22)
    ${\bf y}_{0,j+1}=\frac{{\bf z}_{0,j}}{j},\quad j=0,\ldots,d,$
    and,
    ${\bf y}_{0,0}={\bf y}_{0}-\sum_{k=1}^{M}({\bf y}_{k,0}+{\bf y}_{-k,0}).$
4 Numerical illustrations
In this section we include two numerical examples to illustrate the performance of the Taylor-Fourier approximations proposed in this paper, when compared with state-of-the-art methods. All codes we have used to carry out the numerical experiments are available in https://github.com/jmakazaga/Taylor-Fourier.
4.1 The cubic nonlinear Schrödinger equation
The first example is the cubic nonlinear Schrödinger (NLS) equation
$iu_{t}+u_{xx}+|u|^{2}u=0,\quad x\in[0,2\pi],$ (26)
with periodic boundary conditions and prescribed initial data
$u(x,0)=u^{0}(x).$ (27)
For the numerical solution of (26)-(27), we use the method of lines and then, we first discretize in space with the spectral collocation method as follows.
We fix an integer $J$, set $h=\pi/J$, introduce the uniform grid of points $x_{j}=(j-1)h,1\leq j\leq 2J$, and denote ${\bf U}(t)=[U_{1}(t),\ldots,U_{2J}(t)]^{T}$, where $U_{j}(t)$ is an approximation to $u(x_{j},t)$, for $1\leq j\leq 2J$. We also consider the vector $\hat{\bf U}(t)=[\hat{U}_{1}(t),\ldots,\hat{U}_{2J}(t)]^{T}$ obtained by applying the discrete Fourier transform to ${\bf U}(t)$, that is, for $k=1,\ldots,2J$,
$\hat{U}_{k}(t)=\frac{1}{2J}\sum_{j=1}^{2J}U_{j}(t)e^{-i(k-1)x_{j}}.$
In matrix notation, we write it as $\hat{\bf U}={\cal F}{\bf U}$. The vector ${\bf U}$ can be recovered from $\hat{\bf U}$ by applying the inverse Fourier transform, i.e., ${\bf U}={\cal F}^{-1}\hat{\bf U}$. More explicitly, for $j=1,\ldots,2J$,
$U_{j}(t)=\frac{1}{2J}\sum_{k=1}^{2J}\hat{U}_{k}(t)e^{i(k-1)x_{j}},$
or equivalently,
$U_{j}(t)=\sum_{k=-J}^{J}\hat{u}_{k}(t)e^{ikx_{j}},$
where
$\begin{array}[]{l}\hat{u}_{0}(t)=\dfrac{1}{2J}\hat{U}_{1}(t),\quad\hat{u}_{J}(t)=\hat{u}_{-J}(t)=\dfrac{1}{4J}\hat{U}_{J+1}(t),\\[8.0pt] \hat{u}_{k}(t)=\dfrac{1}{2J}\hat{U}_{k+1}(t),\quad\hat{u}_{-k}(t)=\dfrac{1}{2J}\hat{U}_{2J-k+1}(t),\quad k=1,\ldots,J-1.\end{array}$ (28)
Notice that
$\tilde{u}(x,t)=\sum_{k=-J}^{J}\hat{u}_{k}(t)e^{ikx}$
is a trigonometric polynomial satisfying the interpolation conditions
$\tilde{u}(x_{j},t)=U_{j}(t),\quad j=1,\ldots,2J.$
Both, the discrete Fourier transform and its inverse, can be efficiently computed by applying the fast Fourier transform (FFT) algorithm. This is particularly efficient when $J=2^{l}$.
The spectral collocation semi-discretization of the nonlinear Schrödinger equation is obtained by requiring that
$i\tilde{u}_{t}(x_{j},t)+\tilde{u}_{xx}(x_{j},t)+|\tilde{u}(x_{j},t)|^{2}\tilde{u}(x_{j},t)=0,\quad\quad j=1,\ldots,2J.$
By taking into account that
$i\,\tilde{u}_{xx}(x_{j},t)$ $=$ $\sum_{k=-J}^{J}-i\,k^{2}\,\hat{u}_{k}(t)e^{ikx_{j}}$
$=$ $\frac{1}{2J}\sum_{k=1}^{J}\lambda_{k}\,\hat{U}_{k}(t)e^{i(k-1)x_{j}},$
where
$\lambda_{k}=\left\{\begin{array}[]{rcl}-i(k-1)^{2}&\mbox{ for }&k=1,\ldots,J+1,\\ -i(2J-k+1)^{2}&\mbox{ for }&k=J+2,\ldots,2J,\end{array}\right.$
we arrive at the semi-discretized NLS equation
$\frac{d}{dt}{\bf U}=A{\bf U}+{\bf g}({\bf U}),$ (29)
where
${\bf g}({\bf U})=i\,\left(\begin{matrix}|U_{1}|^{2}U_{1}\\ \vdots\\ |U_{2J}|^{2}U_{2J}\end{matrix}\right),$ (30)
$A={{\cal F}}^{-1}D{{\cal F}}$, and $D$ is the diagonal matrix with diagonal entries
$(\lambda_{1},\ldots,\lambda_{2J})=(0,-i,-2^{2}i,\ldots,-(J-1)^{2}i,-J^{2}i,-(J-1)^{2}i,\ldots,-2^{2}i,-i).$
The time-dependent change of variables ${\bf U}=e^{tA}{\bf W}$ transforms (29) into
$\frac{d}{dt}{\bf W}={\bf f}(t,{\bf W}),$ (31)
where ${\bf f}(t,{\bf W})$ is $(2\pi)$-periodic in $t$ and is defined by
${\bf f}(\theta,{\bf W})=e^{-\theta A}{\bf g}\left(e^{\theta A}{\bf W}\right).$
Products of a vector in $\mathbb{C}^{2J}$ by matrices $e^{\theta A}$ and $e^{-\theta A}$ can be computed efficiently by performing two FFT in $\mathbb{C}^{2J}$, since $e^{\theta A}={\cal F}^{-1}e^{\theta D}{\cal F}$, and $e^{\theta D}$ is the diagonal matrix with diagonal entries $(e^{\lambda_{1}\theta},\ldots,e^{\lambda_{2J}\theta})$. On the other hand, due to the particular form of the nonlinearity ${\bf g}$ in (30), the Taylor expansions required to apply the Taylor-Fourier integrator are cheaply available since sums and products of power series in $t$ can be computed in a very efficient way.
In the numerical experiments we have considered the NLS equation with initial data $u(x,0)=\eta(x)$ equal to the step function
$\eta(x)=\left\{\begin{array}[]{l}-1\quad\mbox{ if }x\in[0,\pi),\\ \hphantom{-}1\quad\mbox{ if }x\in[\pi,2\pi],\end{array}\right.$ (32)
a rough initial data that has widely appeared in the recent literature [13, 14, 6]. For instance, in [6] the initial condition (32) is used to illustrate the Talbot effect.
Figure 1: Numerical solution of (26)-(32) at $t=0.3$ (top left), $t=0.31$ (top right), $t=0.314$ (bottom left) and $t=\pi/10$ (bottom right) computed with $J=2^{9}$, and the $(2^{11},3)$-Taylor-Fourier approximation. Real part is plotted with blue solid line and imaginary part with brown dashed line.
In the first experiment, we take $J=2^{7}$, $2^{8}$, $2^{9}$, $2^{10}$, $T=\pi/10$, $d\leq 5$, and for each $J$, $M=J$, $2J$, $4J$. Once computed the coefficients of the $(M,d)$-Taylor-Fourier approximation, we have evaluated ${\bf W}^{[d]}(t)$ at $t=0.3$, $0.31$, $0.314$ and $\pi/10$, as in [6]. The results for $J=2^{9}$, $M=4J$ and $d=3$ are shown in Figure 1, to be compared with those in Figures 5 and 6 in [6]. Different from [6], we have included in the same plot the real and the imaginary part of the numerical solutions.
Secondly, to illustrate how the choice of $d$ and $M$ affects the numerical solution computed with the Taylor-Fourier method, we include in Figure 2 the numerical approximations to the solution of (26)-(32) at $t=\pi/10$ when $J=2^{9}$, and the combinations $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right) are used. Plots in Figure 2 must be compared with the bottom right plot in Figure 1 where the $(4J,3)$-Taylor-Fourier approximation is displayed. Comparing the plots corresponding to the same value of $M$ and reducing $d$ from 3 to 1, we observe that, although the overall shape of the numerical solutions is rather similar, the maximum and minimum values of both solutions are clearly different. Looking now at the solutions obtained with the same value of $d$ we see that when decreasing $M$ from $4J$ to $J$ small oscillations appear.
Figure 2: $(M,d)$-Taylor-Fourier approximations to the solution of (26)-(32) at $t=\pi/10$ for $J=2^{9}$, computed with $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right). Real part is plotted with blue solid line and imaginary part with brown dashed line.
In order to test the effect of changing the basic frequency $\omega$, we consider (26) with initial data
$u(0,x)=\epsilon\,\eta(x)$ (33)
and $t\in[0,\epsilon^{-2}\pi/10]$, with different values of $\epsilon>0$. This is equivalent to considering $u(t,x)=\epsilon\,v(\epsilon^{2}\,t,x)$, where $v(\tau,x)$ is the solution of
$i\,v_{\tau}=-\omega\,v_{xx}-|v|^{2}v,\quad x\in[0,2\pi],\quad\tau\in[0,\pi/10],$
with $\omega=\epsilon^{-2}$.
In the abstract and in Section 1 we claim that the accuracy of Taylor-Fourier approximations does not deteriorate as the basic frequency $\omega$ increases. We next check whether this holds true in the particular case of the ODE system obtained after the space discretization with $J=2^{6}$ of (26) for the $(2^{11},7)$-Taylor-Fourier approximation, with initial data (33) for different values of $\epsilon>0$. In Figure 3, we represent (in logarithmic scale) the error $|\tilde{U}_{j}(t)-U_{j}(t)|$ ($j=1,\ldots,2J$) at time $t=\epsilon^{-2}\pi/10$ versus $x_{j}$, $j=1,\ldots,2J$ for $\epsilon=2^{-m}$, $m=1,2,3,4$. Here, $\tilde{U}_{j}(t)$ denotes the reference solution of the semi-discretized system computed using the Lawson integrator based on the implementation DP5 in the julia library DifferentialEquation.jl of the 5th order embedded explicit Runge-Kutta pair proposed in [8], with tolerance for absolute and relative errors equal to $10^{-12}$. We observe that the errors do not increase as $\omega=\epsilon^{-2}$ increases, but rather decrease linearly with $\epsilon=\omega^{-1/2}$ in this case.
Figure 3: Errors of the $(2^{11},7)$-Taylor-Fourier approximation at $t=\epsilon^{-2}\pi/10$ for $\epsilon=2^{-m}$, $m=1,2,3,4$ with $J=2^{6}$ and initial data (33).
Finally, we have tested the efficiency (time-discretization error $|\tilde{U}_{j}(t)-U_{j}(t)|$ versus CPU time) of Taylor-Fourier approximations to the solution $u(t,x)$ of (26) with initial data (33). We have considered different space discretization parameters depending on the value of $\epsilon$: $J=2^{9}$ for $\epsilon=1$, and $J=2^{6}$ for $\epsilon=2^{-4}$. We have plotted in Figure 4 the results corresponding to $(M,d)$-Taylor-Fourier integrations with different choices of the parameters $M$ and $d$.
The reference solution of the space-discretized problem used to measure time-discretization errors has been computed as indicated above. For comparison purposes, we have also included errors obtained with the same scheme for tolerances $10^{-3}$, $\ldots$, $10^{-6}$ when $\epsilon=1$, and for tolerances $10^{-3}$, $\ldots$, $10^{-7}$ for $\epsilon=2^{-4}$. In both plots, a black dashed line representing the spatial error for the chosen value of $J$ (estimated by a precise time-integration of the semi-discretized problem with $2J$ spatial nodes) has been included. Notice that although $J$ is larger for $\epsilon=1$ than for $\epsilon=2^{-4}$, the spatial error is also larger for $\epsilon=1$ than for $\epsilon=2^{-4}$. This is due to the factor $\epsilon$ that multiplies the rough initial data (32). In both cases we have used $2\leq d\leq 5$.
Figure 4: Errors against CPU time when $J=2^{9}$, the initial data is (32) and $T=\pi/10$ (left), and when $J=2^{6}$, the initial data is (33) with $\epsilon=2^{-4}$ and $T=\epsilon^{-2}\pi/10$ (right). Diamonds correspond to the Lawson method based on DP5 and circles to the $(M,d)$-Taylor-Fourier approximations
We observe in the left plot of Figure 4 that the Taylor-Fourier integrations are considerably less efficient than the Lawson-DP5 integrations for the case $\epsilon=1$. In the right plot, for $\epsilon=2^{-4}$, we see that the performance of both integrators is rather similar. The low efficiency of the Taylor-Fourier methods for $\epsilon=1$ is due to the high dimension of the semi-discretized problem ($J=2^{9}$), and to the fact that in order to get temporal errors not larger than the spatial errors, the parameter $M$ varies from $2^{8}$ to $2^{11}$, increasing the computational cost of the method. For $\epsilon=2^{-4}$, the dimension of the semi-discretized problem has been divided by $8$, and the problem is more difficult to be solved using standard integrators because its solution is really highly oscillatory. Then, although $M$ ranges between $2^{6}$ and $2^{9}$, the advantage of Lawson-DP5 with respect to Taylor-Fourier is much smaller.
4.2 A perturbed Kepler problem
The second example is a perturbed Kepler problem (the main problem of satellite theory), which accounts for the oblateness of the earth. More precisely, it is the Kepler problem perturbed by the $J_{2}$-term,
$\ddot{\bf q}$ $=-\frac{\mu}{r^{3}}{\bf q}-\nabla V({\bf q}),$ (34)
${\bf q}(t_{0})$ $={\bf q}_{0},\quad\dot{\bf q}(t_{0})=\dot{\bf q}_{0},$ (35)
where
${\bf q}=\left(\begin{matrix}x\\ y\\ z\end{matrix}\right),\quad r=\|{\bf q}\|=\sqrt{x^{2}+y^{2}+z^{2}},\quad\dot{\bf q}=\frac{d}{dt}{\bf q},\quad\ddot{\bf q}=\frac{d^{2}}{dt^{2}}{\bf q},$
and
$V({\bf q})=J_{2}\,\frac{\mu\,R_{e}^{2}}{2\,r^{3}}\,\left(3\,\sin^{2}\theta-1\right),\quad\sin\theta=\frac{z}{r}.$ (36)
Here, $J_{2}$ is the zonal harmonic coefficient ($J_{2}=1.08262668\times 10^{-3}$), $\mu$ is the standard gravitational parameter of the Earth ($\mu=398600.44189\,km^{3}/seg^{2}$) and $R_{e}$ is the equatorial radius of the Earth ($R_{e}=6378.137\,km$).
From now on, we will use the notation $\epsilon:=J_{2}\,\mu\,R_{e}^{2}$. From (36), we get that
$\nabla V(x,y,z)=\frac{3\,\epsilon}{2\,r^{5}}\left(\begin{matrix}(1-5\,\sin^{2}\theta)\,x\\ (1-5\,\sin^{2}\theta)\,y\\ (3-5\,\sin^{2}\theta)\,z\end{matrix}\right).$
4.2.1 KS-formulation of the equations of motion
We will rewrite the equations by making use of the Kustaanheimo-Stiefel (KS) transformation.
The position variables ${\bf q}=(x,y,z)\in\mathbb{R}^{3}$ are replaced by new variables ${\bf u}=(u_{1},u_{2},u_{3},u_{4})\in\mathbb{R}^{4}$ satisfying the relation ${\bf q}=L({\bf u}){\bf u}$, where
$L({\bf u})=\left(\begin{matrix}u_{1}&-u_{2}&-u_{3}&u_{4}\\ u_{2}&u_{1}&-u_{4}&-u_{3}\\ u_{3}&u_{4}&u_{1}&u_{2}\\ \end{matrix}\right).$
Clearly, it holds that $\|{\bf u}\|^{2}=\|{\bf q}\|=r$.
Instead of the physical time $t$, a new independent variable (the fictitious time)
[…]
will be used (we will denote with a prime the derivative with respect to $\tau$), and the physical time $t$ is considered as an additional dependent variable of the equations of motion.
Following [16], the solution curve ${\bf q}(t)$ of the initial value problem (34)–(35) can be parametrized as ${\bf q}=L({\bf u}(\tau)){\bf u}(\tau)$, $t=t(\tau)$, where $({\bf u}(\tau),t(\tau))$ is the solution of the initial value problem
${\bf u}^{\prime\prime}$ $=-\frac{h}{2}{\bf u}-\nabla R({\bf u}),$ (37)
${\bf u}(0)$ $={\bf u}_{0},\quad{\bf u}^{\prime}(0)={\bf u}^{\prime}_{0},$ (38)
$t^{\prime}$ $=\|{\bf u}\|^{2},$ (39)
$t(0)$ $=t_{0},$ (40)
where
$h=\frac{\mu}{\|{\bf q}_{0}\|}-\frac{1}{2}\,\langle\dot{\bf q}_{0},\dot{\bf q}_{0}\rangle-V({\bf q}_{0}),$
and
$R({\bf u})=\frac{1}{2}\,\|{\bf u}\|^{2}\,V(L({\bf u}){\bf u})=\frac{\epsilon}{4\,\|{\bf u}\|^{4}}\,\left(3\,\sin^{2}\theta-1\right),$
with $\sin\theta=2\,(u_{1}u_{3}+u_{2}u_{4})/\|{\bf u}\|^{2}$.
The initial condition ${\bf u}_{0}\in\mathbb{R}^{4}$ is chosen in such a way that
${\bf q}_{0}=L({\bf u}_{0}){\bf u}_{0},$ (41)
and ${\bf u}_{0}^{\prime}\in\mathbb{R}^{4}$ is determined as
${\bf u}_{0}^{\prime}=\frac{1}{2}\,L({\bf u}_{0})^{T}\dot{\bf q}_{0}.$ (42)
There are infinitely many choices of ${\bf u}_{0}\in\mathbb{R}^{4}$ satisfying (41) for a given ${\bf q}_{0}=(x_{0},y_{0},z_{0})$. Among them, we will uniquely determine ${\bf u}_{0}=(u_{1,0},u_{2,0},u_{3,0},u_{4,0})$ by imposing the following conditions:
$\begin{array}[]{ll}\mbox{if }x_{0}\geq 0,\quad u_{1,0}=u_{4,0}\geq 0,\\ \mbox{if }x_{0}<0,\quad u_{2,0}=u_{3,0}\geq 0.\end{array}$ (43)
This leads to
$u_{1,0}=u_{4,0}=\frac{1}{2}\,\sqrt{r_{0}+x_{0}},\quad u_{2,0}=\frac{y_{0}u_{1,0}+z_{0}u_{4,0}}{r_{0}+x_{0}},\quad u_{3,0}=\frac{z_{0}u_{1,0}-y_{0}u_{4,0}}{r_{0}+x_{0}},$
if $x_{0}\geq 0$, and
$u_{2,0}=u_{3,0}=\frac{1}{2}\,\sqrt{r_{0}-x_{0}},\quad u_{1,0}=\frac{y_{0}u_{2,0}+z_{0}u_{3,0}}{r_{0}-x_{0}},\quad u_{4,0}=\frac{z_{0}u_{2,0}-y_{0}u_{3,0}}{r_{0}-x_{0}},$
if $x_{0}<0$, where $r_{0}=\|{\bf q}_{0}\|$.
The gradient of $R({\bf u})$ can be computed as
$\nabla R({\bf u})=\ \frac{\epsilon}{\|{\bf u}\|^{6}}\,(1-6\sin^{2}\theta)\left(\begin{matrix}u_{1}\\ u_{2}\\ u_{3}\\ u_{4}\end{matrix}\right)+\frac{3\,\epsilon}{\|{\bf u}\|^{6}}\,\sin\theta\left(\begin{matrix}u_{3}\\ u_{4}\\ u_{1}\\ u_{2}\end{matrix}\right).$
Clearly, the system of second order ODEs (37) is of the form (1) with
$\omega=\sqrt{h/2},\quad{\bf x}=\left(\begin{matrix}{\bf u}\\ {\bf u}^{\prime}\end{matrix}\right),\quad A=\left(\begin{matrix}0&\omega^{-1}I_{4}\\ -\omega\,I_{4}&0\end{matrix}\right),\quad g({\bf x})=\left(\begin{matrix}{\bf 0}\\ -\nabla R({\bf u})\end{matrix}\right),$
where $I_{4}$ denotes the $4\times 4$ identity matrix.
4.2.2 Stiefel and Scheifele’s formulation
From now on, we assume that $h>0$, and denote $\omega=\sqrt{h/2}$. The solution ${\bf u}(\tau)$ of the initial value problem (37)–(38) can be written as [16]
${\bf u}(\tau)=\cos(\omega\tau)\bm{\alpha}(\tau)+\omega^{-1}\sin(\omega\tau)\bm{\beta}(\tau),$
where $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are the solutions of the 8-dimensional ODE system
$\begin{split}\bm{\alpha}^{\prime}&=\omega^{-1}\sin(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\\ \bm{\beta}^{\prime}&=-\cos(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\end{split}$ (44)
supplemented with the initial conditions
$\bm{\alpha}(0)={\bf u}_{0},\quad\bm{\beta}(0)={\bf u}^{\prime}_{0}.$ (45)
Actually, (44) is the VOP formulation (4) corresponding to (37).
The solution $t(\tau)$ of (39)–(40) can be written as
[…]
which can be explicitly computed once appropriate $(M,d)$-Taylor-Fourier approximations
$\bm{\alpha}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\alpha}_{k,j},\quad\bm{\beta}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\beta}_{k,j},$
of $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are obtained.
4.2.3 High precision numerical propagation of a geostationary orbit
We now consider the initial state of a geostationary satellite [12, pg. 116]. The initial position coordinates (in kilometers) and velocity vectors (in kilometers per second) are
${\bf q}_{0}$ $=(42149.1336,0,0),$
$\dot{\bf q}_{0}$ =(0,3.075823259987749,0.0010736649055318406).absent03.0758232599877490.0010736649055318406=(0,3.075823259987749,0.0010736649055318406).= ( 0 , 3.075823259987749 , 0.0010736649055318406 ) .
Recall from the discussion above that the initial conditions (45) for the equations (44) can be computed from (41), (42) and (43).
We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the VOP formulation (44), and found that $M=8$ and $d=8$ is a good choice in that particular initial value problem. For comparison purposes, we have solved (44) numerically (with absolute and relative tolerance set to $10^{-13}$) for 400 revolutions with an explicit RK method of order 9 due to Verner [18], efficiently implemented with adaptive time-stepping as algorithm Vern9 in DifferentialEquations.jl. We have estimated the errors of both numerical approximations by comparing them with a reference solution computed (in high precision floating point arithmetic) with Vern9 with absolute and relative tolerance set to $10^{-20}$.
Figure 5: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(8,8)$-Taylor-Fourier approximation (bottom).
In Figure 5 we display the relative errors in position ${\bf q}$ (left subplot) and the absolute errors in physical time $t$ (right subplot) versus the fictitious time $\tau$ scaled by the period $P=2\pi/\omega$ of the satellite orbits. The upper (resp. lower) plots correspond to errors of Vern9 (resp. Taylor-Fourier). Observe that the absolute errors in physical time are smaller for the Taylor-Fourier approximation. Moreover, while similar relative errors in position are obtained for both approximations at the end of the integration interval, computing the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of 20 on our laptop).
We emphasize that while Vern9 exhibits the expected pattern of increasing position errors due to the propagation of local truncation and round-off errors, the Taylor-Fourier approximation maintains positional errors below a threshold of approximately $3\times 10^{-15}$ (corresponding to the errors due to truncation of Fourier modes) for approximately 380 revolutions. Using $M=16$ (not shown here) instead of $M=8$, that threshold lowers to round-off levels. Observe that after $380$ revolutions, these errors start to grow. We have checked that error growth is approximately proportional to $t^{d+1}$. The validity of the Taylor-Fourier approximation can be extended to longer time intervals by increasing the degree $d$. For instance, with $d=9$, the errors in position remain below the above-mentioned threshold for up to 500 revolutions.
4.2.4 High precision numerical propagation of a highly eccentric orbit
We now make similar computations with the initial state of a highly eccentric orbit (with eccentricity $e=0.7679436$) obtained from [12, pg. 51]. In particular,
${\bf q}_{0}$ =(11959.886901183693,16289.448826603336,5963.757695165331),absent11959.88690118369316289.4488266033365963.757695165331=(11959.886901183693,-16289.448826603336,-5963.757695165331),= ( 11959.886901183693 , - 16289.448826603336 , - 5963.757695165331 ) ,
$\dot{\bf q}_{0}$ =(4.724300951633136,1.1099935305609756,0.3847854410416176).absent4.7243009516331361.10999353056097560.3847854410416176=(4.724300951633136,-1.1099935305609756,-0.3847854410416176).= ( 4.724300951633136 , - 1.1099935305609756 , - 0.3847854410416176 ) .
Figure 6: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(128,14)$-Taylor-Fourier approximation (bottom).
We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the corresponding initial value problem (44), and found that $M=128$ and $d=14$ gives a precision similar to that obtained with $M=8$ and $d=8$ for the geostationary orbit. This is due to the eccentricity of the current orbit which is much higher than the eccentricity $e\approx 10^{-3}$ of the geostationary orbit, and makes the initial value problem (44) more difficult to be solved numerically (in the sense of requiring more computational effort to achieve similar accuracy). An additional implication of the higher eccentricity is that the interval of validity of the Taylor-Fourier approximations is reduced.
Figure 6 is similar to Figure 5, the only differences being that we display the errors for up to 40 revolutions (instead of 400) and that $(M,d)=(128,14)$ (instead of $(M,d)=(8,8)$). Vern9 is still applied with absolute and relative tolerances set to $10^{-13}$.
As in the previous example, the absolute errors in physical time are smaller for the Taylor-Fourier approximation. In this second example, the relative errors in position are also smaller for the Taylor-Fourier approximation, while the computation of the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of $2.6$ on our laptop).
In this example, the $(128,14)$-Taylor-Fourier approximation maintains position errors below a threshold of approximately $8\times 10^{-16}$ up to $35$ revolutions, and beyond that, exhibits an error growth that is closely proportional to $t^{d+1}$. We have checked that this threshold is not lowered by increasing $M$, unless a higher precision floating point arithmetic is employed.
5 Conclusions
We have proposed approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$. These approximate solutions, that can be written in closed form combining truncated Fourier and Taylor series, are valid for a relatively large time interval and, once computed, can be efficiently evaluated at arbitrary times in that interval. The resulting approximate solutions are uniformly accurate in the basic frequency $\omega$, which makes them more efficient than standard numerical integrators when $\omega$ is large enough. In addition, we have provided a detailed procedure to compute efficiently such approximate solutions making use of power series arithmetic techniques and the FFT algorithm.
We have also reported numerical results to show the performance of the Taylor-Fourier approximations when compared with state-of-the-art methods. When applied to a perturbed Kepler problem, the proposed approximations become more efficient than the numerical integration of the transformed system using the adaptive algorithm Vern9, even for highly eccentric orbits. The application of the Taylor-Fourier integration procedure to the semi-discretized cubic nonlinear Schrödinger equation is not so advantageous from the computational point of view when compared with the Lawson integrator based on the variable-step implementation of DP5. This is mainly due to the high dimension of the semi-discretized problem. However, the uniform accuracy of Taylor-Fourier approximations with respect to the basic frequency $\omega$ implies that they are more efficient than standard numerical methods (which typically experience accuracy deterioration as the frequency $\omega$ increases) for high enough values of $\omega$.
A detailed error analysis of Taylor-Fourier approximations, and their connection with other related topics such as stroboscopic averaging and the solution of the transport equation will be the subject of future work.
Acknowledgements. All the authors have received funding by the Spanish State Research Agency through project PID2022-136585NB-C22,
MCIN/AEI/10.13039/501100011033, European Union. M.P.C. was also supported by project VA169P20 (Junta de Castilla y León, ES) cofinanced by FEDER funds (EU). J.M. and A.M. were partially supported by the Department of Education of the Basque Government through the Consolidated Research Group MATHMODE (ITI456-22).

### v2

@@
-Taylor-Fourier integration
+Taylor-Fourier approximation
 Abstract
-In this paper we introduce an algorithm which provides approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$.
-The proposed approximate solutions are written in closed form as functions $X(t,\omega\,t)$ where $X(t,\theta)$ is, (i) a truncated Fourier series in $\theta$ for fixed $t$, and (ii) a truncated Taylor series in $t$ for fixed $\theta$ (that is the reason for the name of the proposed integrators).
-Such approximations are intended to be uniformly accurate in $\omega$ (in the sense that their accuracy is not deteriorated as $\omega\to\infty$).
-This feature implies that Taylor-Fourier approximations become more efficient than the application of standard numerical integrators for sufficiently high basic frequency $\omega$.
-The main goal of the paper is to propose a procedure to efficiently compute such approximations by combining power series arithmetic techniques and the FFT algorithm.
-We present numerical experiments that demonstrate the effectiveness of our approximation method through its application to well-known problems of interest.
-Keywords: Taylor series, Fourier methods, oscillatory problems, numerical integrators.
+In this paper, we introduce an algorithm that provides approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions, which, after an appropriate change of variables, can be rewritten as non-autonomous systems with a $(2\pi/\omega)$-periodic dependence on $t$.
+The proposed approximate solutions are given in closed form as functions $X(\omega t,t)$, where $X(\theta,t)$ is (i) a truncated Fourier series in $\theta$ for fixed $t$ and (ii) a truncated Taylor series in $t$ for fixed $\theta$, which motivates the name of the method.
+These approximations are uniformly accurate in $\omega$, meaning that their accuracy does not degrade as $\omega\to\infty$.
+In addition, Taylor-Fourier approximations enable the computation of high-order averaging equations for the original semi-linear system, as well as related maps that are particularly useful in the highly oscillatory regime (i.e., for sufficiently large $\omega$).
+The main goal of this paper is to develop an efficient procedure for computing such approximations by combining truncated power series arithmetic with the Fast Fourier Transform (FFT).
+We present numerical experiments that illustrate the effectiveness of the proposed method, including applications to the nonlinear Schrödinger equation with non-smooth initial data and a perturbed Kepler problem from satellite orbit dynamics.
+Keywords: Taylor series, Fourier methods, oscillatory ODE problems, numerical approximations.
 1 Introduction
-In this paper we consider differential systems of the form
+In this paper we are interested in systems of ordinary differential equations of the form
+$\frac{d}{dt}{\bf x}=\mathcal{A}{\bf x}+{\bf g}({\bf x}),$ (1)
+where ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a smooth map and $\mathcal{A}$ is a real $D\times D$ matrix with purely imaginary eigenvalues.
+Although the core concepts and techniques can be naturally extended to the general case of arbitrary purely imaginary eigenvalues, we primarily focus on the fully resonant case, in which the ratio of any pair of eigenvalues is rational.
+In this scenario, the solution ${\bf x}(t)=e^{t\mathcal{A}}{\bf x}(0)$ to (1) with ${\bf g}\equiv{\bf 0}$ becomes $(2\pi/\omega)$-periodic in $t$, where $\omega$ is the fundamental frequency (i.e., all eigenvalues of $\mathcal{A}$ are integer multiples of $\omega i$.)
+We thus assume that $\mathcal{A}=\omega\,A$, where $A$ is a real $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$.
+We focus on approximating the solution ${\bf x}(t)$ of
 $\frac{d}{dt}{\bf x}=\omega A{\bf x}+{\bf g}({\bf x}),$ (2)
-where ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a smooth map, $\omega$ is a positive real parameter, and $A$ is a real $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$.
-Hence, the matrix exponential $e^{\theta A}$ is $(2\pi)$-periodic in $\theta$ and, therefore, the solution ${\bf x}(t)=e^{t\omega A}{\bf x}(0)$ of (1) with ${\bf g}\equiv{\bf 0}$ is $(2\pi/\omega)$-periodic in $t$.
-If $\omega$ is large, the solution of the homogeneous problem associated to (1) is highly oscillatory.
+supplemented with the initial condition
+${\bf x}(0)={\bf x}_{0}\in\mathbb{R}^{D}.$ (3)
 Examples of differential systems (2) with these properties can be found, for instance, when dealing with semi-discretized dispersive partial differential equations [3, 11, 19], or in some formulations of the equations of motion in satellite dynamics [18].
-We focus on approximating the solution ${\bf x}(t)$ of (1) supplemented with the initial condition
-${\bf x}(0)={\bf x}_{0}\in\mathbb{R}^{D}.$ (2)
+In the present work, we aim at obtaining a closed form approximation to the solution ${\bf x}(t)$ of (2) that exhibits uniform accuracy across varying frequencies $\omega$ for a time-interval (centered at $t=0$) of length independent of $\omega$.
 With the change of variables
@@
 the initial value problem (2)–(3) reads
-$\frac{d}{dt}{\bf y}={\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf x}_{0},$ (4)
+$\frac{d}{dt}{\bf y}={\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf y}_{0},$ (5)
 with
 ${\bf f}(\theta,{\bf y})=e^{-\theta A}{\bf g}(e^{\theta A}{\bf y}),$ (6)
-which is $(2\pi)$-periodic in $\theta$.
-Provided that the basic frequency $\omega$ is high enough, solving numerically (4) by a Runge-Kutta method demonstrates superior efficiency compared to the direct application of the same Runge-Kutta scheme to the original equation (1).
+which is $(2\pi)$-periodic in $\theta$, and ${\bf y}_{0}={\bf x}_{0}$.
 In the context of problems of the form (2) obtained from the space semi-discretization of dispersive PDEs, the numerical integration with a Runge-Kutta method of the transformed system (5), as a means to solve (2)–(3), leads to the so-called Lawson’s generalized Runge-Kutta methods [13, 11].
@@
 Other VOP formulations (obtained by using different sets of coordinates) of the equations of motion in satellite dynamics are also of the form (5).
-From now on we will focus on computing approximations to the solution ${\bf y}(t)$ of the initial value problem (4) which are valid on an appropriate interval around $t=0$.
-In the highly oscillatory case, for high enough basic frequency $\omega$, the larger is the basic frequency $\omega$, the more computational effort will be required by the Runge-Kutta method to achieve a prescribed accuracy.
-In the present work, we aim at obtaining a closed form approximation to the solution ${\bf y}(t)$ of (4) that exhibits uniform accuracy across varying frequencies $\omega$.
-This implies that comparable accuracy is attained with the same computational effort even for arbitrarily high frequencies.
-The Taylor-Fourier approximations that we propose here draw inspiration from previous works on the numerical integration of oscillatory problems (see e.g.
-[4], [7], and references therein), and are related to the methods proposed in [5] and in [2].
-For a prescribed positive integer $M$, we propose to compute a sequence
-$\left\{{\bf y}^{[1]}(t),{\bf y}^{[2]}(t),\ldots,{\bf y}^{[d]}(t),\ldots\right\}$
-of approximations to the solution ${\bf y}(t)$ of (4) of the form
-${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right),\quad d=1,2,\ldots,$ (6)
-with appropriate coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, and starting from ${\bf y}^{[0]}(t)\equiv{\bf x}_{0}$ (this means that we can use (6) with $d=0$, setting ${\bf y}_{0,0}^{[0]}={\bf x}_{0}$).
-We will refer to (6) as the (M,d)-Taylor-Fourier approximation to the solution of (4).
-In order to compute the successive approximations (6) to the solution of (4), for $d=0,1,\ldots$, we do the following:
-  • Given ${\bf y}^{[d]}(t)$, we first insert it into the right hand side of (4), and then we determine an approximation ${\bf z}^{[d]}(t)$ to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ of the form
-    ${\bf z}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf z}_{k,j}^{[d]}\right).$ (7)
-    This approximation, to be defined precisely in Section 2, is determined by combining truncated Taylor expansions and trigonometric interpolation of periodic functions.
-That is the reason for the name of the proposed integrators.
-  • Once ${\bf z}^{[d]}(t)$ has been obtained, we use it as the right-hand side of (4) and integrate to obtain
+The Taylor-Fourier approximations that we propose here draw inspiration from previous works on the analysis of numerical integration of oscillatory problems:
+  • The technique of modulated Fourier expansions  (see e.g. [8, 10], and references therein), when applied to (5) would lead to a formal expansion
+    $\sum_{k\in\mathbb{Z}}e^{ik\omega t}{\bf y}_{k}(t)$ (7)
+    of the solution ${\bf y}(t)$ of (5), where ${\bf y}_{k}(t)$ ($k\in\mathbb{Z}$) are smoothly varying functions (i.e., their derivatives are bounded independently of $\omega$).
+An approximation of ${\bf y}(t)$ can be obtained from (7) by replacing each ${\bf y}_{k}(t)$ by its $d$th degree truncated Taylor expansion centered at $t=0$, and by discarding the terms corresponding to Fourier modes with $|k|>M$, that is,
+    ${\bf y}(t)\approx\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}{\bf y}_{k,j}.$ (8)
+    The $(M,d)$-Taylor-Fourier approximations that we propose in the present paper are precisely of the form (8).
+  • In [5], the high-order averaging of systems of the form (5) is studied by expanding its solution ${\bf y}(t)$ as a formal series indexed by multi-indices $(k_{1}\cdots k_{\ell})\in\mathbb{Z}^{\ell}$.
+Such series expansion can be best described by parametrizing the equations (5) with a dummy variable $\epsilon$ as
+    $\frac{d}{dt}{\bf y}=\epsilon\,{\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf y}_{0}.$ (9)
+    Its solution ${\bf y}(t)$ is expanded as
+    ${\bf y}(t)={\bf y}_{0}+\sum_{\ell=1}^{\infty}\epsilon^{\ell}\sum_{(k_{1}\cdots k_{\ell})\in\mathbb{Z}^{\ell}}\alpha_{k_{1}\cdots k_{\ell}}(t)\,{\bf f}_{k_{1}\cdots k_{\ell}}({\bf y}_{0}),$
+    where each ${\bf f}_{k_{1}\cdots k_{\ell}}$ is a map that is independent of $\omega$, and each $\alpha_{k_{1}\cdots k_{\ell}}(t)$ is a scalar function that only depends on $t$ and $\omega$ and the multi-index $(k_{1},\ldots,k_{\ell})$, but is independent from the actual map ${\bf f}(\theta,{\bf y})$.
+More precisely, each $\alpha_{k_{1}\cdots k_{\ell}}(t)$ is a linear combination with rational coefficients of terms of the form $\omega^{j-\ell}t^{j}e^{i\,k\,\omega\,t}$, $0\leq j\leq\ell$.
+An equivalent expansion with series indexed by mode-colored rooted trees was derived in [4].
+After collecting terms with equal powers of $\epsilon$ such expansions of the solution ${\bf y}(t)$ can be written as
+    ${\bf y}(t)={\bf y}_{0}+\sum_{\ell=1}^{\infty}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf y}_{k,\ell,j}.$ (10)
+    As shown in [5, Theorem 3.1], under suitable conditions on the map ${\bf f}(\theta,{\bf y})$, there exists $\eta>0$ independent of $\omega$ such that the series (10) is absolutely and uniformly convergent for $|\epsilon\,t|\leq\eta$.
+An essential insight that inspired the consideration of Taylor-Fourier approximations, as presented in this paper, is the realization [4] that the approximations derived from the expansion (10) by truncating terms involving higher powers of $\epsilon$,
+    ${\bf y}^{[d]}(t)={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf y}_{k,\ell,j},$ (11)
+    can be systematically constructed by combining Picard iterations with the truncation of the series expansion in powers of $\epsilon$.
+Indeed, starting with ${\bf y}^{[0]}(t)={\bf y}_{0}$, for $d=1,2,3,\ldots$, one first determines
+    ${\bf z}^{[d]}(t)=\sum_{\ell=0}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf z}_{k,\ell,j}$ (12)
+    uniquely such that
+    ${\bf f}(\omega t,{\bf y}^{[d]}(t))={\bf z}^{[d]}(t)+\mathcal{O}(\epsilon^{d+1}),$ (13)
+    and then obtains
 […]
-    It is straightforward to check that, by construction, ${\bf y}_{k,d+1}^{[d+1]}=0$ for $k\neq 0$.
-The aim of the present paper is: (i) to introduce and give a precise definition of Taylor-Fourier approximations, (ii) to propose a procedure to compute them efficiently by making extensive use of the FFT algorithm, and (iii) to illustrate the application of the proposed approximation method to certain well known problems of interest.
-Our findings suggest that, provided that the basic frequency $\omega$ is high enough, Taylor-Fourier approximations can achieve comparable precision to standard numerical integration methods applied to the transformed system (4), while demanding less CPU time.
-We want to stress that employing Taylor-Fourier approximations may be of interest even in cases where the application of standard numerical integrators to the transformed system (4) proves to be very efficient.
-Indeed, the algorithm introduced in the present paper provides, for given $d$ and $M$, closed expressions
-${\bf X}(t)=e^{t\omega A}\left(\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right)$ (9)
-of approximate solutions of (1)-(2), valid for a relatively large time interval and that, once computed, can be efficiently evaluated at arbitrary times $t$.
-This is in contrast with most time integrators commonly used, which generate approximations to the exact solution of the problem either at a prescribed final time $T$ or at a finite set $0<t_{1}<t_{2}<\cdots\leq T$ of intermediate times.
-(In the later case, approximations to the solutions at arbitrary times $t\in[0,T]$ are typically obtained with piecewise interpolating polynomials.)
-In our numerical experiments, we will compare the $(M,d)$-Taylor-Fourier approximations (6) to the solution ${\bf y}(t)$ of (4) (equivalently, the approximation (9) to the solution ${\bf x}(t)$ of (1)–(2)) with the approximations obtained by numerically integrating (4) with highly efficient Runge-Kutta methods (equivalently, numerically integrating (1)–(2) with Lawson’s generalized Runge-Kutta methods).
-A state-of-the-art highly performant library for the numerical integration of differential equations, DifferentialEquation.jl [15], implemented in julia language [1], is freely available and can be used to integrate the transformed system (4) with optimized adaptive explicit Runge-Kutta integrators.
-In order to compare such numerical solvers with the integrators proposed in this paper, we have carried out a careful implementation of the Taylor-Fourier methods in julia language, paying special attention to computational aspects such as memory management and efficient power series arithmetic.
-In Section 2 we give a precise definition of $(M,d)$-Taylor-Fourier approximations, and demonstrate their applicability to problem classes extending beyond real semi-linear problems of the form (1).
-Section 3 is devoted to describe an algorithm designed to efficiently compute Taylor-Fourier approximations.
-In Section 4 we present a couple of relevant numerical examples where the performance of Taylor-Fourier integrators is analyzed.
-Finally, in Section 5, we summarize our conclusions and outline avenues for future research on Taylor-Fourier methods.
-2 Taylor-Fourier methods
-2.1 Definition of Taylor-Fourier approximations
-So far, we have not specified how to determine the approximation (7) to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ in the procedure introduced in Section 1.
-We first note that (6) and (7) can be rewritten as
-${\bf y}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\omega t),\qquad{\bf z}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\omega t),$
-respectively, where each component of the $D$-vector functions ${\bf Y}_{j}^{[d]}(\theta)$ and ${\bf Z}_{j}^{[d]}(\theta)$ is a trigonometric polynomial of degree $M$ in the angular variable $\theta$, that is,
-${\bf Y}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf y}_{k,j}^{[d]},\qquad{\bf Z}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j}^{[d]},$
-with coefficients ${\bf y}_{k,j}^{[d]},{\bf z}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$.
-Given a positive integer $M$ and ${\bf Y}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$, we will determine ${\bf Z}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$ by imposing the following conditions:
-  • For each $\theta_{n}=n\pi/M$, $n=0,1,2,\ldots,2M-1$,
-    ${\bf f}\left(\theta_{n},\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta_{n})\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta_{n})+\mathcal{O}(t^{d+1}){\color[rgb]{1,0,0},}$ (10)
-    as $t\to 0$.
-  • For $j=0,1,\ldots,d$,
-    ${\bf z}_{-M,j}^{[d]}={\bf z}_{M,j}^{[d]}.$ (11)
-The first condition is aimed at ensuring that
-$\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta)\approx{\bf f}\left(\theta,\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)\right)$
-for all $\theta\in[0,2\pi]$ and for sufficiently small $t\in\mathbb{R}$, whereas the second one is included to ensure uniqueness.
-To sum up, given a positive integer $M$, the successive approximations (6) to the solution of (4) are uniquely determined by (10)–(11) and (8).
-We will refer to the approximation (6) as the $(M,d)$-Taylor-Fourier approximation.
-2.2 Application to more general equations
-The definition of Taylor-Fourier approximations can be extended without modifications to initial value problems of the form (4) in the complex domain.
-In particular, Taylor-Fourier approximations can be successfully applied to equations of the form (1) defined in the complex domain, provided that $A$ is a complex $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$, and the map ${\bf g}:\mathbb{C}^{D}\to\mathbb{C}^{D}$ is such that ${\bf g}({\bf x})={\bf G}({\bf x},{\bf\bar{x}})$, where ${\bf G}:\mathbb{C}^{2D}\to\mathbb{C}^{D}$ is a polynomial map.
-Going back to equations in the real domain, Taylor-Fourier methods are not limited to semi-linear equations of the form (1). They can be applied more generally to problems of the form
-$\frac{d}{dt}{\bf x}=\omega\,{\bf r}({\bf x})+{\bf g}({\bf x}),\quad{\bf x}(0)={\bf x}_{0},$ (12)
-where ${\bf r}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ and ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ are smooth maps and $\omega$ is a positive real parameter, provided that the following two assumptions hold:
+    In the particular case where $\epsilon=1$, one obtains a series expansion of the solution ${\bf y}(t)$ of (5) that is absolutely and uniformly convergent for $|t|\leq\eta$, where $\eta>0$ is independent of $\omega$.
+    Taylor-Fourier approximations, as presented in this paper, emerged as a practical computational method for approximating such expansions.
+To construct approximations of the form (8), an iterative process similar to (12)–(14) is employed.
+However, instead of truncating the series in powers of $\epsilon$, truncation is applied to powers of $t$, and the exact Fourier series expansion is replaced with trigonometric interpolation.
+Our Taylor-Fourier approximations are also related to the methods proposed in [2] for a more restricted class of problems of the form (2).
+The aim of the present paper is: (i) to introduce and give a precise definition of Taylor-Fourier approximations, (ii) to relate them to high order stroboscopic averaging, (iii) to propose a procedure to compute them efficiently by making use of truncated power series arithmetic and the FFT algorithm, and (iv) to illustrate the application of the proposed approximation method to certain well known problems of interest.
+For given $M$ and $d$, the algorithm introduced in this paper provides closed-form expressions
+${\bf x}(t)\approx e^{t\omega A}\left(\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right)$ (15)
+of approximate solutions to (2)-(3).
+These approximations maintain uniform accuracy in $\omega$ and remain valid over a time interval independent of $\omega$.
+Once computed, these approximations can be efficiently evaluated at arbitrary times $t$ without additional costly computations.
+This contrasts with standard numerical integration methods, which produce approximations to the exact solution at a finite sequence of intermediate times $0<t_{1}<t_{2}<\cdots\leq T$, and whose accuracy typically deteriorates for large values of $\omega$.
+The Taylor-Fourier approximations presented in this work are not intended to compete with numerical integration methods in terms of computational efficiency.
+Rather, their primary purpose is to provide reliable reference solutions over large time intervals.
+Additionally, they can serve as a valuable tool for assessing the accuracy and effectiveness of methods that approximate the solutions of high-order averaged equations.
+We have carried out a careful implementation of the Taylor-Fourier methods in Julia language, paying special attention to computational aspects such as memory management and efficient truncated power series arithmetic.
+In Section 2 we give a precise definition of $(M,d)$-Taylor-Fourier approximations.
+Section 3 is devoted to describe an algorithm designed to efficiently compute truncated Taylor-Fourier series.
+In Section 4, we explore the relationship of the Taylor-Fourier technique with high-order stroboscopic averaging.
+We demonstrate in Section 5 the applicability of Taylor-Fourier approximations to problem classes extending beyond real semi-linear problems of the form (1) with fully resonant frequencies.
+In Section 6 we illustrate the application of the Taylor-Fourier algorithm to a couple of relevant examples.
+Finally, in Section 7, we summarize our conclusions.
+2 Taylor-Fourier approximations
+To lay the groundwork for the definition of fully discrete $(M,d)$-Taylor-Fourier approximations (8) to the solution ${\bf y}(t)$ of (5) for arbitrary pairs of positive integers $(M,d)$, we first introduce the semi-discrete $(\infty,d)$-Taylor-Fourier approximations, which correspond to the limit case as $M\to\infty$.
+2.1 Semi-discrete Taylor-Fourier approximations
+For each positive integer $d$, the $(\infty,d)$-Taylor-Fourier approximation to the solution ${\bf y}(t)$ of (5) is of the form
+${\bf Y}^{[d]}(\omega t,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\omega t),$ (16)
+where each component of the $D$-vector functions ${\bf Y}_{j}^{[d]}(\theta)$ is $(2\pi)$-periodic in the angular variable $\theta$. The function
+${\bf Y}^{[d]}(\theta,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)$
+is defined recursively. Starting with ${\bf Y}^{[0]}(\theta,t)={\bf y}_{0}$, ${\bf Y}^{[d+1]}(\theta,t)$ is determined for $d\geq 0$ as follows:
   1. 1.
-    For each $t\in\mathbb{R}$ there exists a $t$-flow map $\varphi_{t}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ of the differential system $\frac{d}{dt}{\bf x}={\bf r}({\bf x})$.
-That is, for each $t\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{D}$,
-    $\frac{\partial}{\partial t}\varphi_{t}({\bf x})={\bf r}(\varphi_{t}({\bf x})){\color[rgb]{1,0,0}.}$ (13)
-    In that case, problem (12) is transformed with the change of variables
-    ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$
-    into (4), where
-    ${\bf f}(\theta,{\bf y})=\varphi^{\prime}_{\theta}({\bf y})^{-1}{\bf g}(\varphi_{\theta}({\bf y})).$ (14)
-    Here, $\varphi^{\prime}_{\theta}({\bf y})$ denotes the Jacobian matrix of the map $\varphi_{\theta}$ evaluated at the vector ${\bf y}$.
+    First, we determine the $d$th-order truncated expansion in powers of $t$ for ${\bf f}(\theta,{\bf Y}^{[d]}(\theta,t))$,
+    ${\bf Z}^{[d]}(\theta,t)=\sum_{j=0}^{d}t^{j}{\bf Z}_{j}^{[d]}(\theta)={\bf f}\left(\theta,\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)\right)+\mathcal{O}(t^{d+1}).$ (17)
   2. 2.
-    The function ${\bf f}(\theta,{\bf y})$ defined in (14) is $(2\pi)$-periodic in $\theta$.
-Under these two assumptions, the solution ${\bf x}(t)$ of (12) can be approximated by
-${\bf x}^{[d]}(t)=\varphi_{\omega\,t}\left({\bf y}^{[d]}(t)\right),$ (15)
-where
-${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}$
-is the $(M,d)$-Taylor-Fourier approximation to the solution of (4)–(14).
-In particular, the above-mentioned two assumptions hold for systems of the form
-$\begin{split}\frac{d}{dt}\phi_{j}&=k_{j}\omega+g_{j}({\bf x}),\quad j=1,\ldots,d,\\ \frac{d}{dt}a_{j}&=g_{d+j}({\bf x}),\quad j=1,\ldots,D-d,\\ \end{split}$ (16)
-where ${\bf x}=(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})$, $k_{j}\in\mathbb{Z}$, and for each $j$, $g_{j}$ is $2\pi$-periodic in each of the angle variables $\phi_{j}$. In that case,
-$\varphi_{t}(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})=\left(\begin{matrix}\phi_{1}+k_{1}\,t\\ \vdots\\ \phi_{d}+k_{d}\,t\\ a_{1}\\ \vdots\\ a_{D-d}\\ \end{matrix}\right),$
-and the change of variables ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$, with ${\bf y}=(\tilde{\phi}_{1},\ldots,\tilde{\phi}_{d},\tilde{a}_{1},\ldots,\tilde{a}_{D-d})$, transforms the system (16) into
-$\frac{d}{dt}\tilde{\phi}_{j}$ $=g_{j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,d,$
-$\frac{d}{dt}\tilde{a}_{j}$ $=g_{d+j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,D-d.$
-3 Efficient implementation of Taylor-Fourier integrators
-We will next show, for given positive integers $M$ and $d$, and given vector coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$, how to compute efficiently the new coefficients ${\bf y}_{k,j}^{[d+1]}\in\mathbb{C}^{D}$, $10\leq j\leq d+$ 1, $-M\leq k\leq M$.
-This computation is carried out in four sequential steps (for simplicity of notation, we omit superscripts $[d]$ in ${\bf Y}_{j}^{[d]}$ and ${\bf Z}_{j}^{[d]}$, when they are not necessary).
-We recall that $\theta_{n}=n\pi/M,\quad n=0,1,\ldots,2M-1$.
+    Second, we determine ${\bf Y}^{[d]}(\theta,t)$ from the identity
+[…]
+Notice that if, for each $j$, the function ${\bf Y}_{j}^{[d]}(\theta)$ admits a Fourier expansion
+${\bf Y}_{j}^{[d]}(\theta)=\sum_{k=-\infty}^{\infty}e^{ik\theta}{\bf y}^{[d]}_{k,j},$
+then the $(\infty,d)$-Taylor-Fourier approximation
+${\bf Y}^{[d]}(\omega t,t)=\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}$ (19)
+takes the form (8) in the limiting case $M\to\infty$.
+2.2 Fully discrete Taylor-Fourier approximations
+For the time being, let us assume that the Fourier expansion of ${\bf f}(\theta,{\bf y})$ contains only a finite number of nonzero modes. That is, there exists a positive integer $M_{0}$ such that
+${\bf f}(\theta,{\bf y})=\sum_{k=-M_{0}}^{M_{0}}e^{ik\theta}\,\hat{\bf f}_{k}({\bf y}).$ (20)
+In particular, if each component of ${\bf g}({\bf x})$ in (2) is a polynomial function of ${\bf x}\in\mathbb{R}^{D}$, then (6) admits such a finite Fourier expansion.
+If condition (20) holds, then the semi-discrete Taylor-Fourier approximation (19) can also be expressed as a finite sum.
+More precisely, it can be shown that ${\bf y}^{[d]}_{k,j}=0$ for all $|k|>M=dM_{0}$.
+In this case, the first step (17) in the recursive definition of $(\infty,d)$-Taylor-Fourier approximations can be split into two sub-steps.
+In the first sub-step, truncated expansions in powers of $t$ of univariate functions of $t$ are performed, while in the second one, we apply trigonometric interpolation for uniformly spaced nodes in $[0,2\pi]$.
+More precisely, we can consider $\theta_{n}=2n\pi/(2M+1)$, $n=0,1,2,\ldots,2M$, with $M\geq(d+1)M_{0}$, and
+  • first compute for each node $\theta_{n}$ the vectors ${\bf Z}^{[d]}_{j}(\theta_{n})={\bf Z}_{n,j}^{[d]}\in\mathbb{C}^{D}$ from the truncated series expansion in powers of $t$
+    ${\bf f}\left(\theta_{n},\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta_{n})\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}^{[d]}+\mathcal{O}(t^{d+1})$ (21)
+  • and then, for each $j=0,1,2,\ldots,d$, uniquely determine the functions
+    ${\bf Z}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j}^{[d]}$ (22)
+    from the conditions
+    ${\bf Z}_{j}^{[d]}(\theta_{n})={\bf Z}_{n,j}^{[d]},\quad\theta_{n}=\frac{2n\pi}{2M+1},\quad n=0,1,2,\ldots,2M.$
+The expansion (21) can be computed by evaluating the left-hand side of (21) with truncated power series arithmetic.
+As for the trigonometric interpolation, it is efficiently performed with the application of the Discrete Fourier Transform (DFT) (see Section 3).
+Since the efficiency of DFT-based trigonometric interpolation is optimal for number of nodes that are powers of 2, interpolation with an even number of nodes is preferred in practice.
+We thus replace the interpolatory conditions (having an odd number of nodes) above by the alternative conditions with an even number of nodes,
+${\bf Z}_{j}^{[d]}(\theta_{n})={\bf Z}_{n,j}^{[d]},\quad\theta_{n}=\frac{n\pi}{M},\quad n=0,1,2,\ldots,2M-1,$ (23)
+together with
+${\bf z}_{-M,j}^{[d]}={\bf z}_{M,j}^{[d]},\quad j=0,1,\ldots,d.$ (24)
+Condition (24) is imposed to ensure uniqueness of the trigonometric interpolation functions (22).
+To sum up, (21)–(24) is a practical way of performing the first step (17) in the recursive definition of $(\infty,d)$-Taylor-Fourier approximations to the solution ${\bf y}(t)$ for (5) provided that the finiteness condition (20) and $M>(d+1)M_{0}$ hold.
+Otherwise, the functions ${\bf Z}^{[d]}_{j}(\theta)$ determined by (21)–(24) do not satisfy (17) exactly, as they will be affected by interpolation errors, that will depend on the actual value of the positive integer $M$.
+In that general scenario, the function
+${\bf y}^{[M,d]}(t)={\bf Y}^{[M,d]}(\omega t,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[M,d]}(\omega t),$ (25)
+defined recursively by (21)–(24) and (18) (with the superscript $[d]$ replaced by $[M,d]$) can be seen as a fully discrete substitute of the $(\infty,d)$-Taylor-Fourier approximation.
+We will say that (25) is the $(M,d)$-Taylor-Fourier approximation to the solution ${\bf y}(t)$ for (5).
+3 Efficient computation of Taylor-Fourier approximations
+Next, we describe an efficient method to compute the new coefficients ${\bf y}_{k,j}^{[M,d+1]}\in\mathbb{C}^{D}$ for $10\leq j\leq d+$ 1 and $-M\leq k\leq M$, given the coefficients ${\bf y}_{k,j}^{[M,d]}\in\mathbb{C}^{D}$ for $0\leq j\leq d$ and $-M\leq k\leq M$.
+For simplicity, we omit the superscripts $[M,d]$ in the remainder of this section.
+This computation proceeds in four sequential steps.
 3.1 Step 1: Evaluation of trigonometric polynomials
 For $n=0,1,\ldots,2M-1$, we calculate the vector coefficients
-${\bf Y}_{n,j}={\bf Y}_{j}(\theta_{n})=\sum_{k=-M}^{M}e^{ikn\pi/M}{\bf y}_{k,j}^{[d]},\quad j=0,1,\ldots,d,$ (17)
-by applying the inverse DFT
+${\bf Y}_{n,j}={\bf Y}_{j}(\theta_{n})=\sum_{k=-M}^{M}e^{ikn\pi/M}{\bf y}_{k,j},\quad j=0,1,\ldots,d,$ (26)
+(with the nodes $\theta_{n}$ considered in (23)) by applying the inverse DFT
 ${\bf Y}_{n,j}=\frac{1}{2M}\sum_{k=0}^{2M-1}e^{ikn\pi/M}\hat{\bf Y}_{k,j},\quad n=0,1,\ldots,2M-1,$
 where for $j=0,1,\ldots,d$,
-$\begin{array}[]{l}\hat{\bf Y}_{0,j}=2M\,{\bf y}_{0,j}^{[d]},\quad\hat{\bf Y}_{M,J}=2M({\bf y}_{M,j}^{[d]}+{\bf y}_{-M,j}^{[d]}),\\[8.0pt] \hat{\bf Y}_{k,j}=2M\,{\bf y}_{k,j}^{[d]},\quad\hat{\bf Y}_{2M-k,j}=2M{\bf y}_{-k,j}^{[d]},\quad k=1,\ldots,M-1.\end{array}$
+$\begin{array}[]{l}\hat{\bf Y}_{0,j}=2M\,{\bf y}_{0,j},\quad\hat{\bf Y}_{M,j}=2M({\bf y}_{M,j}+{\bf y}_{-M,j}),\\[8.0pt] \hat{\bf Y}_{k,j}=2M\,{\bf y}_{k,j},\quad\hat{\bf Y}_{2M-k,j}=2M{\bf y}_{-k,j},\quad k=1,\ldots,M-1.\end{array}$
 3.2 Step 2: Truncated Taylor expansions
 For each $n=0,1,\ldots,2M-1$, we compute ${\bf Z}_{n,j}$, $j=0,1,\ldots,d$, such that
-${\bf f}\left({\theta_{n},}\sum_{j=0}^{d}t^{j}\,{\bf Y}_{n,j}\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}+{\cal O}(t^{d+1}).$
-This means that we obtain the Taylor polynomial of degree $d$ of the left hand side of the previous equality.
-The computation of the required Taylor expansion can be done in a very efficient way when appropriate tools to perform power series arithmetic are available.
+${\bf f}\left({\theta_{n},}\sum_{j=0}^{d}t^{j}\,{\bf Y}_{n,j}\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}+{\cal O}(t^{d+1}).$ (27)
+The required truncated Taylor expansion (27) can be computed efficiently without resorting to symbolic calculations by evaluating the univariate function on the left-hand side of (27) using truncated power series arithmetic.
+To achieve this, one can implement the mapping ${\bf f}(\theta,{\bf y})$ as a function that takes an angle $\theta$ and $D$ polynomials in $t$ of degree $d$, returning the corresponding $D$ polynomials of degree $d$.
+This approach has been used in both examples presented in Section 6.
+Alternatively, numerous libraries are available across different computing environments that support arithmetic for truncated power series.
 3.3 Step 3: Trigonometric interpolation
@@
 ${\bf z}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf z}_{k,j}\right),$
-we want to write
+we want to compute
 […]
-To simplify the notation, we will drop the superscript $[d+1]$ from the vector coefficients ${\bf y}_{k,j}^{[d+1]}$ in the remaining of the present subsection, so that we will write
-${\bf y}^{[d+1]}(t)=t^{d+1}\,{\bf y}_{0,d+1}+\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}\right).$
 First, we notice that ${\bf z}(t)$ can be written as
@@
     ${\bf y}_{0,0}={\bf y}_{0}-\sum_{k=1}^{M}({\bf y}_{k,0}+{\bf y}_{-k,0}).$
+4 Taylor-Fourier methods and high order stroboscopic averaging
+In this section, we establish a connection between semi-discrete Taylor-Fourier expansions and high-order stroboscopic averaging of systems of the form (9) based on the results in [4].
+We begin by recalling from [4, 5] the series expansion (10) and its relation to high-order averaging, which, as shown in [4], is derived through successive applications of the Picard-like iteration (12)–(14).
+Note that, by rescaling time, we may assume without loss of generality that either $\omega=1$ or $\epsilon=1$ in (9).
+However, we retain this redundancy in this section, as it facilitates the extension of our results to the quasi-periodic case considered in Subsection 5.2.
+In $d$th-order averaging, a change of variables of the form
+${\bf y}={\bf U}(\omega t,{\bf w};\epsilon)={\bf w}+\epsilon\,{\bf u}_{1}(\omega t,{\bf w})+\cdots+\epsilon^{d}\,{\bf u}_{d}(\omega t,{\bf w})$
+is introduced. This transformation converts the original system (9) into an averaged system of the form
+$\frac{d}{dt}{\bf w}=\epsilon\,{\bf F}({\bf w};\epsilon)+\epsilon^{d+1}{\bf R}(\omega t,{\bf w};\epsilon),\quad{\bf w}(0)={\bf w}_{0},$
+where ${\bf F}({\bf w};\epsilon)$ represents the averaged dynamics, and ${\bf R}(\omega t,{\bf w};\epsilon)$ captures the remaining small amplitude oscillatory terms.
+In the stroboscopic averaging framework, the solution ${\bf y}(t)$ of (9) admits the representation
+${\bf y}(t)={\bf U}(\omega t,{\bf W}(t,{\bf y}_{0};\epsilon);\epsilon)+\mathcal{O}(\epsilon^{d+1}),$ (36)
+where ${\bf W}(t,{\bf y}_{0};\epsilon)={\bf w}(t)$ denotes the solution of the averaged system
+$\frac{d}{dt}{\bf w}=\epsilon\,{\bf F}({\bf w};\epsilon),\quad{\bf w}(0)={\bf y}_{0}.$ (37)
+Consequently, the solution ${\bf x}(t)$ of the original semi-linear equation
+$\frac{d}{dt}{\bf x}=\mathcal{A}{\bf x}+\epsilon\,{\bf g}({\bf x})$ (38)
+satisfies
+${\bf x}(t)=e^{t\mathcal{A}}\,{\bf U}(\omega t,{\bf W}(t,{\bf y}_{0};\epsilon);\epsilon)+\mathcal{O}(\epsilon^{d+1}).$
+As shown in [4, Theorem 2.10] (see also [5, Theorem 2.1]), we have that, for arbitrary ${\bf y}_{0}\in\mathbb{R}^{D}$,
+$\begin{split}{\bf U}(\omega t,{\bf y}_{0};\epsilon)&={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf y}_{k,\ell,0},\\ {\bf W}(t,{\bf y}_{0};\epsilon)&={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{j=0}^{\ell}t^{j}\sum_{k=-\infty}^{\infty}{\bf y}_{k,\ell,j},\\ {\bf F}({\bf y}_{0};\epsilon)&=\frac{\partial}{\partial t}{\bf W}(0,{\bf y}_{0};\epsilon)=\sum_{\ell=1}^{d}\epsilon^{\ell-1}\sum_{k=-\infty}^{\infty}{\bf y}_{k,\ell,1},\end{split}$ (39)
+where ${\bf y}_{k,\ell,j}$ are the coefficients of the expansion (10) of the solution of (9).
+Recall that these coefficients can be obtained through successive applications of the Picard-like iteration (12)–(14).
+The $(\infty,d)$-Taylor-Fourier approximation of the solutions ${\bf y}(t)$ and ${\bf x}(t)$ of (9) and (38) read
+${\bf Y}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})=\sum_{j=0}^{d}\epsilon^{j}t^{j}\,\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf y}_{k,j}^{[d]}.$ (40)
+and ${\bf X}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})=e^{t\,\mathcal{A}}\,{\bf Y}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})$, respectively.
+Expansions (40) and (11) are derived through a similar Picard-like iteration process.
+The key distinction lies in the truncation method used to approximate ${\bf z}(t)={\bf f}(\omega t,{\bf y}(t))$: (40) is obtained by truncating higher powers of $t$, whereas (11) involves truncating higher powers of $\epsilon$.
+This latter approach discards more terms in the expansion of ${\bf z}^{[d]}(t)={\bf f}(\omega t,{\bf y}^{[d]}(t))$.
+This implies that the coefficients of the expansions (40) and (11) are related by
+$\epsilon^{j}{\bf y}_{k,j}^{[d]}=\epsilon^{j}{\bf y}_{k,j,j}+\mathcal{O}(\epsilon^{d+1}).$
+We thus have the following relationships:
+${\bf U}(\omega t,{\bf y}_{0};\epsilon)$ $={\bf Y}^{[d]}(\omega t,0;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d+1}),$
+${\bf W}(t,{\bf y}_{0};\epsilon)$ $={\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d+1}),$
+$\epsilon\,{\bf F}({\bf y}_{0};\epsilon)$ $=\left.\frac{\partial}{\partial t}{\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})\right|_{t=0}+\mathcal{O}(\epsilon^{d+1}).$
+This demonstrates that the change of variables in the $d$th-order stroboscopic averaging can be effectively approximated by applying the $(M,d)$-Taylor-Fourier algorithm with a sufficiently large $M$.
+In particular, this together with (36) implies that
+${\bf X}^{[d]}\left(\omega t,0;\epsilon,{\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})\right)={\bf X}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d}).$ (41)
+While the right-hand side of the averaged equations and their solutions can also be computed using this approach, we do not claim it to be an efficient method for this purpose.
+However, it can serve as a useful tool to assess the accuracy of more efficient implementations of approximate averaged equations and their solutions.
+We note that the semi-discrete expansions (11) in powers of $\epsilon$ can be transformed into a practical fully discrete method by adapting the procedure proposed in this paper.
+Specifically, for a given pair $(M,d)$, a finite-sum substitute for the $d$th order expansion (11) can be constructed by applying trigonometric interpolation to replace the periodic functions
+$\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf z}_{k,\ell,j}$
+in the first sub-step of the iteration defined by (12)–(14).
+This approach is particularly relevant in the context of stroboscopic averaging, especially when one aims to evaluate the mappings in (39) exactly –or with high precision for a sufficiently large $M$– for a given order $d$.
+In particular, the right-hand side ${\bf F}({\bf y}_{0};\epsilon)$ of the averaged equations obtained through this method preserves the geometric structure of the original equations.
+For instance, if the original equations (38) are Hamiltonian, then the averaged equations retain this Hamiltonian structure.
+In contrast, the approximation
+$\left.\frac{\partial}{\partial t}{\bf Y}^{[M,d]}(0,t;\epsilon,{\bf y}0)\right|_{t=0}\approx{\bf F}({\bf y}_{0};\epsilon),$
+obtained using the Taylor-Fourier algorithm proposed in Section 3, generally does not correspond to a Hamiltonian vector field.
+In this context, the truncated expansions (11) and (40) are not symplectic, as is expected for any explicit formula derived by truncating a power series expansion of a symplectic map.
+However, this fully discrete method is computationally more expensive than the Taylor-Fourier algorithm proposed in this paper, both in terms of memory and complexity.
+For a given order $d$, it requires $(d+2)/2$ times more coefficients, and the algorithm involves arithmetic with truncated power series in two variables (rather than one).
+5 Application to more general equations
+5.1 Semi-linear equations in the complex domain
+The definition of Taylor-Fourier approximations can be extended without modifications to initial value problems of the form (5) in the complex domain.
+In particular, Taylor-Fourier approximations can be successfully applied to equations of the form (2) defined in the complex domain, provided that $A$ is a complex $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$, and the map ${\bf g}:\mathbb{C}^{D}\to\mathbb{C}^{D}$ is such that ${\bf g}({\bf x})={\bf G}({\bf x},{\bf\bar{x}})$, where ${\bf G}:\mathbb{C}^{2D}\to\mathbb{C}^{D}$ is a polynomial map.
+5.2 Semi-linear equations with non-resonant frequencies
+Taylor-Fourier approximations can be naturally extended to systems of the form (1), where $\mathcal{A}$ is a real $D\times D$ matrix with purely imaginary eigenvalues that are not fully resonant.
+Applying the time-dependent change of variables
+${\bf x}(t)=e^{t\mathcal{A}}{\bf y}(t),$ (42)
+we transform (1) into
+$\frac{d}{dt}{\bf y}=e^{-t\mathcal{A}}{\bf g}(e^{t\mathcal{A}}{\bf y}).$
+In this general setting, there exists a vector of non-resonant frequencies $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{r})\in\mathbb{R}^{r}$ such that $\mathbf{k}\cdot\boldsymbol{\omega}=k_{1}\omega_{1}+\cdots+k_{r}\omega_{r}\neq 0$ for all $\mathbf{k}=(k_{1},\ldots,k_{r})\in\mathbb{Z}^{r}/\{{\bf 0}\}$.
+The eigenvalues of $\mathcal{A}$ then take the form $(\mathbf{k}\cdot\boldsymbol{\omega})i$ for $\mathbf{k}\in\mathbb{Z}^{r}$, making the solution operator $e^{t\mathcal{A}}$ of the linear part of (1) quasi-periodic in $t$, rather than strictly periodic as in the fully resonant case.
+This implies that there exists a mapping
+$\begin{split}{\bf f}:&\mathbb{T}^{r}\times\mathbb{R}^{D}\longrightarrow\mathbb{R}^{D},\\ &\ (\boldsymbol{\theta},{\bf y})\mapsto{\bf f}(\boldsymbol{\theta},{\bf y})\end{split}$
+such that
+${\bf f}(\boldsymbol{\omega}t,{\bf y})\equiv e^{-t\mathcal{A}}{\bf g}(e^{t\mathcal{A}}{\bf y}).$
+Therefore, the change of variables (42) transforms (1) into the quasi-periodic system
+$\frac{d}{dt}{\bf y}={\bf f}(\boldsymbol{\omega}t,{\bf y}).$ (43)
+The definition of semi-discrete Taylor-Fourier approximations follows naturally: in Subsection 2.1, replace the basic frequency $\omega$ and the angular variable $\theta$ by the vector of non-resonant frequencies $\boldsymbol{\omega}$ and the angular variables $\boldsymbol{\theta}$, respectively.
+The discussion on high order stroboscopic averaging in Section 4 is also valid in the quasi-periodic case with the referred changes (the averaging technique is referred in that case as quasi-stroboscopic averaging [4, 5]).
+Fully discrete $(M,d)$-Taylor-Fourier approximations extend this approach by employing trigonometric interpolation of functions $\mathbb{T}^{r}\to\mathbb{C}$ over an $r$-fold Cartesian product of the uniform grid
+$\theta_{n}=\frac{n\pi}{M},\quad n=0,1,2,\ldots,2M-1,$
+in $[0,2\pi)$, which can be efficiently computed using multi-dimensional discrete Fourier transforms (DFT).
+However, compared to the periodic case, quasi-periodicity requires increased memory and higher computational cost: while a complex tensor of dimensions $D\times(d+1)\times(M+1)$ suffices to store the coefficients ${\bf y}^{[M,d]}_{k,j}$ in the periodic case, the quasi-periodic setting with $\boldsymbol{\omega}\in\mathbb{R}^{r}$ requires a complex tensor of dimensions $D\times(d+1)\times(M+1)^{r}$, significantly increasing storage and processing demands.
+5.3 Equations in other alternative formats
+Going back to equations in the real domain, Taylor-Fourier approximations are not limited to semi-linear equations of the form (2). They can be applied more generally to problems of the form
+$\frac{d}{dt}{\bf x}=\omega\,{\bf r}({\bf x})+{\bf g}({\bf x}),\quad{\bf x}(0)={\bf x}_{0},$ (44)
+where ${\bf r}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ and ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ are smooth maps and $\omega$ is a positive real parameter, provided that the following two assumptions hold:
+  1. 1.
+    For each $t\in\mathbb{R}$ there exists a $t$-flow map $\varphi_{t}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ of the differential system $\frac{d}{dt}{\bf x}={\bf r}({\bf x})$.
+That is, for each $t\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{D}$,
+    $\frac{\partial}{\partial t}\varphi_{t}({\bf x})={\bf r}(\varphi_{t}({\bf x})).$ (45)
+    In that case, problem (44) is transformed with the change of variables
+    ${\bf x}(t)=\varphi_{\omega t}({\bf y}(t))$
+    into (5), where
+    ${\bf f}(\theta,{\bf y})=\varphi^{\prime}_{\theta}({\bf y})^{-1}{\bf g}(\varphi_{\theta}({\bf y})).$ (46)
+    Here, $\varphi^{\prime}_{\theta}({\bf y})$ denotes the Jacobian matrix of the map $\varphi_{\theta}$ evaluated at the vector ${\bf y}$.
+  2. 2.
+    The function ${\bf f}(\theta,{\bf y})$ defined in (46) is $(2\pi)$-periodic in $\theta$.
+Under these two assumptions, the solution ${\bf x}(t)$ of (44) can be approximated by
+${\bf x}^{[M,d]}(t)=\varphi_{\omega\,t}\left({\bf y}^{[M,d]}(t)\right),$ (47)
+where
+${\bf y}^{[M,d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[M,d]}$
+is the $(M,d)$-Taylor-Fourier approximation to the solution of (5) with the map ${\bf f}$ defined by (46).
+In particular, the above-mentioned two assumptions hold for systems of the form
+$\begin{split}\frac{d}{dt}\phi_{j}&=k_{j}\omega+g_{j}({\bf x}),\quad j=1,\ldots,s,\\ \frac{d}{dt}a_{j}&=g_{s+j}({\bf x}),\quad j=1,\ldots,D-s,\\ \end{split}$ (48)
+where ${\bf x}=(\phi_{1},\ldots,\phi_{s},a_{1},\ldots,a_{D-s})$, $k_{j}\in\mathbb{Z}$, and for each $j$, $g_{j}$ is $2\pi$-periodic in each of the angle variables $\phi_{j}$. In that case,
+$\varphi_{t}(\phi_{1},\ldots,\phi_{s},a_{1},\ldots,a_{D-s})=\left(\begin{matrix}\phi_{1}+k_{1}\,t\\ \vdots\\ \phi_{s}+k_{s}\,t\\ a_{1}\\ \vdots\\ a_{D-s}\\ \end{matrix}\right),$
+and the change of variables ${\bf x}(t)=\varphi_{\omega t}({\bf y}(t))$, with ${\bf y}=(\tilde{\phi}_{1},\ldots,\tilde{\phi}_{s},\tilde{a}_{1},\ldots,\tilde{a}_{D-s})$, transforms the system (48) into
+$\frac{d}{dt}\tilde{\phi}_{j}$ $=g_{j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{s}+k_{s}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-s}),\quad j=1,\ldots,s,$
+$\frac{d}{dt}\tilde{a}_{j}$ $=g_{s+j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{s}+k_{s}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-s}),\quad j=1,\ldots,D-s.$
 6 Numerical illustrations
-In this section we include two numerical examples to illustrate the performance of the Taylor-Fourier approximations proposed in this paper, when compared with state-of-the-art methods.
+In this section we include two numerical examples to illustrate the performance of the Taylor-Fourier approximations proposed in this paper.
 All codes we have used to carry out the numerical experiments are available in https://github.com/jmakazaga/Taylor-Fourier.
@@
 Products of a vector in $\mathbb{C}^{2J}$ by matrices $e^{\theta A}$ and $e^{-\theta A}$ can be computed efficiently by performing two FFT in $\mathbb{C}^{2J}$, since $e^{\theta A}={\cal F}^{-1}e^{\theta D}{\cal F}$, and $e^{\theta D}$ is the diagonal matrix with diagonal entries $(e^{\lambda_{1}\theta},\ldots,e^{\lambda_{2J}\theta})$.
-On the other hand, due to the particular form of the nonlinearity ${\bf g}$ in (30), the Taylor expansions required to apply the Taylor-Fourier integrator are cheaply available since sums and products of power series in $t$ can be computed in a very efficient way.
+On the other hand, due to the particular form of the nonlinearity ${\bf g}$ in (53), the Taylor expansions required to compute the Taylor-Fourier approximations are cheaply available since sums and products of power series in $t$ can be computed in a very efficient way.
 In the numerical experiments we have considered the NLS equation with initial data $u(x,0)=\eta(x)$ equal to the step function
@@
 Different from [7], we have included in the same plot the real and the imaginary part of the numerical solutions.
-Secondly, to illustrate how the choice of $d$ and $M$ affects the numerical solution computed with the Taylor-Fourier method, we include in Figure 2 the numerical approximations to the solution of (26)-(32) at $t=\pi/10$ when $J=2^{9}$, and the combinations $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right) are used.
+Secondly, to illustrate how the choice of $d$ and $M$ affects the approximate solution computed with the Taylor-Fourier method, we include in Figure 2 the numerical approximations to the solution of (49)-(55) at $t=\pi/10$ when $J=2^{9}$, and the combinations $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right) are used.
 Plots in Figure 2 must be compared with the bottom right plot in Figure 1 where the $(4J,3)$-Taylor-Fourier approximation is displayed.
@@
 Figure 3: Errors of the $(2^{11},7)$-Taylor-Fourier approximation at $t=\epsilon^{-2}\pi/10$ for $\epsilon=2^{-m}$, $m=1,2,3,4$ with $J=2^{6}$ and initial data (56).
-Finally, we have tested the efficiency (time-discretization error $|\tilde{U}_{j}(t)-U_{j}(t)|$ versus CPU time) of Taylor-Fourier approximations to the solution $u(t,x)$ of (26) with initial data (33).
-We have considered different space discretization parameters depending on the value of $\epsilon$: $J=2^{9}$ for $\epsilon=1$, and $J=2^{6}$ for $\epsilon=2^{-4}$.
-We have plotted in Figure 4 the results corresponding to $(M,d)$-Taylor-Fourier integrations with different choices of the parameters $M$ and $d$.
-The reference solution of the space-discretized problem used to measure time-discretization errors has been computed as indicated above.
-For comparison purposes, we have also included errors obtained with the same scheme for tolerances $10^{-3}$, $\ldots$, $10^{-6}$ when $\epsilon=1$, and for tolerances $10^{-3}$, $\ldots$, $10^{-7}$ for $\epsilon=2^{-4}$.
-In both plots, a black dashed line representing the spatial error for the chosen value of $J$ (estimated by a precise time-integration of the semi-discretized problem with $2J$ spatial nodes) has been included.
-Notice that although $J$ is larger for $\epsilon=1$ than for $\epsilon=2^{-4}$, the spatial error is also larger for $\epsilon=1$ than for $\epsilon=2^{-4}$.
-This is due to the factor $\epsilon$ that multiplies the rough initial data (32).
-In both cases we have used $2\leq d\leq 5$.
-Figure 4: Errors against CPU time when $J=2^{9}$, the initial data is (32) and $T=\pi/10$ (left), and when $J=2^{6}$, the initial data is (33) with $\epsilon=2^{-4}$ and $T=\epsilon^{-2}\pi/10$ (right).
-Diamonds correspond to the Lawson method based on DP5 and circles to the $(M,d)$-Taylor-Fourier approximations
-We observe in the left plot of Figure 4 that the Taylor-Fourier integrations are considerably less efficient than the Lawson-DP5 integrations for the case $\epsilon=1$.
-In the right plot, for $\epsilon=2^{-4}$, we see that the performance of both integrators is rather similar.
-The low efficiency of the Taylor-Fourier methods for $\epsilon=1$ is due to the high dimension of the semi-discretized problem ($J=2^{9}$), and to the fact that in order to get temporal errors not larger than the spatial errors, the parameter $M$ varies from $2^{8}$ to $2^{11}$, increasing the computational cost of the method.
-For $\epsilon=2^{-4}$, the dimension of the semi-discretized problem has been divided by $8$, and the problem is more difficult to be solved using standard integrators because its solution is really highly oscillatory.
-Then, although $M$ ranges between $2^{6}$ and $2^{9}$, the advantage of Lawson-DP5 with respect to Taylor-Fourier is much smaller.
 6.2 A perturbed Kepler problem
@@
 From now on, we assume that $h>0$, and denote $\omega=\sqrt{h/2}$. The solution ${\bf u}(\tau)$ of the initial value problem (60)–(61) can be written as [18]
-${\bf u}(\tau)=\cos(\omega\tau)\bm{\alpha}(\tau)+\omega^{-1}\sin(\omega\tau)\bm{\beta}(\tau),$
-where $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are the solutions of the 8-dimensional ODE system
-$\begin{split}\bm{\alpha}^{\prime}&=\omega^{-1}\sin(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\\ \bm{\beta}^{\prime}&=-\cos(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\end{split}$ (44)
+${\bf u}(\tau)=\cos(\omega\tau)\boldsymbol{\alpha}(\tau)+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}(\tau),$
+where $\boldsymbol{\alpha}(\tau)$ and $\boldsymbol{\beta}(\tau)$ are the solutions of the 8-dimensional ODE system
+$\begin{split}\boldsymbol{\alpha}^{\prime}&=\omega^{-1}\sin(\omega\tau)\nabla R(\cos(\omega\tau)\boldsymbol{\alpha}+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}),\\ \boldsymbol{\beta}^{\prime}&=-\cos(\omega\tau)\nabla R(\cos(\omega\tau)\boldsymbol{\alpha}+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}),\end{split}$ (67)
 supplemented with the initial conditions
-$\bm{\alpha}(0)={\bf u}_{0},\quad\bm{\beta}(0)={\bf u}^{\prime}_{0}.$ (45)
+$\boldsymbol{\alpha}(0)={\bf u}_{0},\quad\boldsymbol{\beta}(0)={\bf u}^{\prime}_{0}.$ (68)
 Actually, (67) is the VOP formulation (5) corresponding to (60).
@@
 which can be explicitly computed once appropriate $(M,d)$-Taylor-Fourier approximations
-$\bm{\alpha}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\alpha}_{k,j},\quad\bm{\beta}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\beta}_{k,j},$
-of $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are obtained.
+$\boldsymbol{\alpha}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\boldsymbol{\alpha}_{k,j},\quad\boldsymbol{\beta}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\boldsymbol{\beta}_{k,j},$
+of $\boldsymbol{\alpha}(\tau)$ and $\boldsymbol{\beta}(\tau)$ are obtained.
 6.2.3 High precision numerical propagation of a geostationary orbit
-We now consider the initial state of a geostationary satellite [12, pg. 116]. The initial position coordinates (in kilometers) and velocity vectors (in kilometers per second) are
+We now consider the initial state of a geostationary satellite [14, pg. 116]. The initial position coordinates (in kilometers) and velocity vectors (in kilometers per second) are
 ${\bf q}_{0}$ $=(42149.1336,0,0),$
@@
 Recall from the discussion above that the initial conditions (68) for the equations (67) can be computed from (64), (65) and (66).
-We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the VOP formulation (44), and found that $M=8$ and $d=8$ is a good choice in that particular initial value problem.
-For comparison purposes, we have solved (44) numerically (with absolute and relative tolerance set to $10^{-13}$) for 400 revolutions with an explicit RK method of order 9 due to Verner [18], efficiently implemented with adaptive time-stepping as algorithm Vern9 in DifferentialEquations.jl.
-We have estimated the errors of both numerical approximations by comparing them with a reference solution computed (in high precision floating point arithmetic) with Vern9 with absolute and relative tolerance set to $10^{-20}$.
-Figure 5: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(8,8)$-Taylor-Fourier approximation (bottom).
-In Figure 5 we display the relative errors in position ${\bf q}$ (left subplot) and the absolute errors in physical time $t$ (right subplot) versus the fictitious time $\tau$ scaled by the period $P=2\pi/\omega$ of the satellite orbits.
-The upper (resp.
-lower) plots correspond to errors of Vern9 (resp.
-Taylor-Fourier).
-Observe that the absolute errors in physical time are smaller for the Taylor-Fourier approximation.
-Moreover, while similar relative errors in position are obtained for both approximations at the end of the integration interval, computing the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of 20 on our laptop).
-We emphasize that while Vern9 exhibits the expected pattern of increasing position errors due to the propagation of local truncation and round-off errors, the Taylor-Fourier approximation maintains positional errors below a threshold of approximately $3\times 10^{-15}$ (corresponding to the errors due to truncation of Fourier modes) for approximately 380 revolutions.
+We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the VOP formulation (67) over 400 revolutions, and found that $M=8$ and $d=8$ is a good choice in that particular initial value problem.
+We have estimated the errors of the Taylor-Fourier approximations by comparing them with a reference solution computed (in high precision floating point arithmetic) with Vern9, an explicit RK method of order 9 due to Verner [20], efficiently implemented with adaptive time-stepping in DifferentialEquations.jl.
+The absolute and relative tolerances have been set equal to $10^{-20}$.
+Figure 4: Relative errors in position (top) and absolute errors in physical time (bottom) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the $(8,8)$-Taylor-Fourier approximation.
+In Figure 4 we display the relative errors in position ${\bf q}$ (top subplot) and the absolute errors in physical time $t$ (bottom subplot) versus the fictitious time $\tau$ scaled by the period $P=2\pi/\omega$ of the satellite orbits for the $(8,8)$-Taylor-Fourier approximation.
+We observe that the Taylor-Fourier approximation maintains positional errors below a threshold of approximately $3\times 10^{-15}$ (corresponding to the errors due to truncation of Fourier modes) for approximately 380 revolutions.
 Using $M=16$ (not shown here) instead of $M=8$, that threshold lowers to round-off levels.
@@
 For instance, with $d=9$, the errors in position remain below the above-mentioned threshold for up to 500 revolutions.
+Figure 5: Energy error of the averaging approximated solution for $(M,d)=(8,8)$ (top) and the position error of the averaged approximation with respect the $(8,8)$-Taylor-Fourier approximation (bottom) in 600 revolutions.
+We next check the accuracy of averaged approximation on the left-hand side of (41).
+We show in Figure 5 the energy error of the averaged approximated solution (top subplot) and the relative error of the position with respect to the Taylor-Fourier approximation (bottom).
+It can be observed that the behavior of the averaged approximation is practically identical to the behavior of the Taylor-Fourier approximation: it maintains the energy error at rounding error levels for approximately 380 revolutions, just like the Taylor-Fourier approximation, and on the other hand, the difference between the averaged approximation and the Taylor-Fourier approximation remains below the rounding error level for more than 600 revolutions.
 6.2.4 High precision numerical propagation of a highly eccentric orbit
-We now make similar computations with the initial state of a highly eccentric orbit (with eccentricity $e=0.7679436$) obtained from [12, pg. 51]. In particular,
+We now make similar computations with the initial state of a highly eccentric orbit (with eccentricity $e=0.7679436$) obtained from [14, pg. 51]. In particular,
 ${\bf q}_{0}$ =(11959.886901183693,16289.448826603336,5963.757695165331),absent11959.88690118369316289.4488266033365963.757695165331=(11959.886901183693,-16289.448826603336,-5963.757695165331),= ( 11959.886901183693 , - 16289.448826603336 , - 5963.757695165331 ) ,
 $\dot{\bf q}_{0}$ =(4.724300951633136,1.1099935305609756,0.3847854410416176).absent4.7243009516331361.10999353056097560.3847854410416176=(4.724300951633136,-1.1099935305609756,-0.3847854410416176).= ( 4.724300951633136 , - 1.1099935305609756 , - 0.3847854410416176 ) .
-Figure 6: Relative errors in position (left) and absolute errors in physical time (right) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the numerical integration with Vern9 with tol=$10^{-13}$ (top) and the $(128,14)$-Taylor-Fourier approximation (bottom).
 We have computed several $(M,d)$-Taylor-Fourier approximations to the solution of the corresponding initial value problem (67), and found that $M=128$ and $d=14$ gives a precision similar to that obtained with $M=8$ and $d=8$ for the geostationary orbit.
@@
 An additional implication of the higher eccentricity is that the interval of validity of the Taylor-Fourier approximations is reduced.
+Figure 6: Relative errors in position (top) and absolute errors in physical time (bottom) versus fictitious time $\tau$ (scaled by $P=2\pi/\omega$) of the $(128,14)$-Taylor-Fourier approximation.
 Figure 6 is similar to Figure 4, the only differences being that we display the errors for up to 40 revolutions (instead of 400) and that $(M,d)=(128,14)$ (instead of $(M,d)=(8,8)$).
-Vern9 is still applied with absolute and relative tolerances set to $10^{-13}$.
-As in the previous example, the absolute errors in physical time are smaller for the Taylor-Fourier approximation.
-In this second example, the relative errors in position are also smaller for the Taylor-Fourier approximation, while the computation of the Taylor-Fourier approximation required considerably less CPU time than Vern9 (by a factor of $2.6$ on our laptop).
 In this example, the $(128,14)$-Taylor-Fourier approximation maintains position errors below a threshold of approximately $8\times 10^{-16}$ up to $35$ revolutions, and beyond that, exhibits an error growth that is closely proportional to $t^{d+1}$.
@@
 We have proposed approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$.
-These approximate solutions, that can be written in closed form combining truncated Fourier and Taylor series, are valid for a relatively large time interval and, once computed, can be efficiently evaluated at arbitrary times in that interval.
-The resulting approximate solutions are uniformly accurate in the basic frequency $\omega$, which makes them more efficient than standard numerical integrators when $\omega$ is large enough.
-In addition, we have provided a detailed procedure to compute efficiently such approximate solutions making use of power series arithmetic techniques and the FFT algorithm.
-We have also reported numerical results to show the performance of the Taylor-Fourier approximations when compared with state-of-the-art methods.
-When applied to a perturbed Kepler problem, the proposed approximations become more efficient than the numerical integration of the transformed system using the adaptive algorithm Vern9, even for highly eccentric orbits.
-The application of the Taylor-Fourier integration procedure to the semi-discretized cubic nonlinear Schrödinger equation is not so advantageous from the computational point of view when compared with the Lawson integrator based on the variable-step implementation of DP5.
-This is mainly due to the high dimension of the semi-discretized problem.
-However, the uniform accuracy of Taylor-Fourier approximations with respect to the basic frequency $\omega$ implies that they are more efficient than standard numerical methods (which typically experience accuracy deterioration as the frequency $\omega$ increases) for high enough values of $\omega$.
-A detailed error analysis of Taylor-Fourier approximations, and their connection with other related topics such as stroboscopic averaging and the solution of the transport equation will be the subject of future work.
+These approximate solutions, that can be written in closed form combining truncated Fourier and Taylor series, are valid over a time interval whose length is independent of $\omega$.
+Once computed, they can be efficiently evaluated at arbitrary times within that interval.
+The resulting approximate solutions are uniformly accurate [6] with respect to the basic frequency $\omega$.
+We have provided a detailed procedure to compute efficiently such approximate solutions making use of power series arithmetic techniques and the FFT algorithm.
+Additionally, we have demonstrated how Taylor-Fourier approximations can be used to evaluate various maps associated with high-order averaging of highly oscillatory systems.
+Furthermore, we have extended the Taylor-Fourier approximation methodology to highly oscillatory ODEs with multiple non-resonant frequencies.
+We have also reported numerical experiments to illustrate the application of the Taylor-Fourier approximations.
+Their uniform accuracy with respect to the basic frequency $\omega$ suggests that they could be more efficient than standard numerical methods (which typically suffer from accuracy deterioration as $\omega$ increases) for sufficiently high values of $\omega$.
+When applied to a perturbed Kepler problem (the so-called $J_{2}$-problem in orbit propagation), the proposed approximations efficiently provide very accurate solutions, even for highly eccentric orbits.
+The Taylor-Fourier approximations to the solution of the semi-discretized cubic nonlinear Schrödinger equation reproduce the well-known Talbot phenomenon, though at the cost of increased computational effort due to the high dimensionality of the semi-discretized problem.
 Acknowledgements. All the authors have received funding by the Spanish State Research Agency through project PID2022-136585NB-C22,